### v1

# Canonical Electrodynamics in Ashtekar-Barbero variables

Abstract
We present the full Hamiltonian analysis of electrodynamics, including fermions, photons, and their interaction, on a general curved spacetime in Ashtekar-Barbero variables. No gauge fixing is performed and the theory is background independent. The details of the constraints of the theory for both Weyl and Dirac fermions are presented. Moreover, the equations of motion are derived and it is shown that they reduce to the equations of motion of the standard QFT in curved or flat spacetime in the appropriate limits. Finally, the behavior of the theory under parity transformation is analyzed.
1 Introduction
Fermions comprise the realistic matter in the universe. Understanding many astrophysical phenomena, particularly those that are relevant for tests of gravity in strong or quantum regimes, need a careful and detailed study of fermions, particularly in the electromagnetic sector, in those regimes. Therefore, it is of utmost importance to have a general background independent formulation of fermions, i.e., one that includes all the backreactions.
Having a classical canonical theory in hand which describes the interaction of these three ingredients is a first and crucial step in being able to proceed with canonical quantization in approaches such as loop quantum gravity (LQG) and generalized uncertainty principle (GUP). Furthermore, there are several interesting symmetry reduced scenarios, such as dynamical (shell collapse, Oppenheimer-Snyder scenario) black hole solutions in the effective regime of LQG, where only scalar fields have been incorporated. Since these models are almost all canonical systems, a readily Hamiltonian formulation where fermions and photons are also involved will open new possibilities for extending these effective models to more realistic scenarios.
There are previous works done with regard to fermions in a canonical theory in Ashtekar-Barbero variables Bojowald:2010qpa ; Das:2008r ; Thiemann:1998QSDV ; Jacobson:1988qta ; Mercuri:2006wb ; PhysRevD.73.084016 ; Date:2011rd that paved the way for our work. However, they usually ignore the interaction term between the fermions and the photons and hence do not present the full QED Hamiltonian in a non-fixed general curved background. Some of them also lack the details of computation or seemingly suffer from some missing terms or some inconsistencies. In particular, it seems that in Das:2008r the choice of the definition of the spin connection $\Gamma_{a}^{i}$ is not compatible with their covariant derivative $D_{a}=\partial_{a}+A_{a}^{i}\tau_{i}$ and their constraints. In Thiemann:1998QSDV , the authors present the Hamiltonian constraint for gravity plus fermions, however, it seems that they do not start directly from the action, and as a consequence there is a missing term in the modification of the Ashtekar connection as well as in the fermionic Hamiltonian constraint. They also seem to take the modification of the Ashtekar connection derived in Jacobson:1988qta . However, the latter starts from the Palatini action rather than the Holst action . As a consequence it does not seem to involve axial fermionic current and only incoporates fermionic current. It also seems that the fact that the extrinsic curvature in presence of fermions is no longer symmetric because of the torsion contribution is overlooked. As a consequence, in the scalar constraint, a term proportional to $\epsilon_{i}{}^{jk}K_{a}^{i}E_{j}^{a}J_{k}$ is missing.
In this work we set to fill the aforementioned gaps. We carefully study a generic background-independent system of gravity, fermions and photons, starting from the action and making a $3+1$ decomposition, writing the theory in Ashtekar-Barbero variables. This is done for both Weyl and Dirac fermions, in a step by step, clear, and detailed manner. We then derive the modification to both the canonical variables, and the constraints, showing in a clear way the distinct contribution of each sector of to these constraints. We also provide the equations of motion of fermions in this setting and study the behavior of the system under parity transformation. Another difference of this work compared to the previous ones, is that we do not modify the gamma matrices, and rather change the identities between them to incorporate the changes needed for the background-independent framework. This way it is easier to go back to a flat QFT theory. Furthermore, as mentioned before, to our knowledge, this is the first time that the QED interaction term is studied in detail in such a canonical formulation in Ashtekar-Barbero variables.
The structure of this paper is as follows: In Sec. 2, we provide a brief overview of vacuum canonical gravity in Ashtekar-Barbero variables starting from the Holst action. Then in Sec. 3.1, we present how Weyl fermions are incorporated in the theory, and how they modify the canonical variables and constraints. In Sec. 3.2 a similar analysis is provided but now for Dirac fermions. Using the results of the previous sections, in Sec. 4.1 the full QED Hamiltonian in an arbitrary dynamical background is constructed by incorporating both the photon field and its interaction with fermions and gravity. In Sec. 4.2 we study the behavior of the system under the parity transformations. In Sec. 4.3 we briefly discuss how the constraints of the gravity-fermion system would modify in case of non-minimal coupling. Finally, in Sec. 5 we summarize and conclude the paper.
2 Brief overview of Ashtekar-Barbero variables
In this section we will quickly review Ashtekar-Barbero formalism of GR in order to introduce the notation used in the Fermions and QED Sections.
We start with a first order action of GR, the Holst action Bojowald:2010qpa
$S=\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}F_{ab}^{KL}(\omega)\,,$ (1)
written in terms of the spacetime tetrads $e_{I}^{a}$ and their determinant $(\text{det}\,e)^{-1}$. The (co)tetrads obey
$g_{ab}=\eta_{IJ}e_{a}^{I}e_{b}^{J}.$ (2)
Throughout this work the first four lower case Latin indices $a,\,b,\,c,\,d$ are abstract spacetime (sometimes spatial) indices. The rest of the lower case Latin indices are $su(2)$ indices. Upper case Latin indices are the internal Lorentz indices associated to the $SO(1,3)$ gauge freedom under which the tetrads transform. The connection one form associated to this transformation is $\omega_{a}^{IJ}$, and $F^{IJ}(\omega)=d\omega^{IJ}+\omega^{I}{}_{K}\wedge\omega^{KJ}$ is the antisymmetric 2-curvature of $\omega_{a}^{IJ}$. Furthermore $\kappa=8\pi G$ and
$P_{KL}^{IJ}=\delta_{K}^{[I}\delta_{L}^{J]}-\frac{1}{2\beta}\epsilon^{IJ}{}_{KL},$ (3)
where $\beta$ is the Barbero-Immirzi parameter. The term in the action corresponding to the second term on the right hand side of $P_{KL}^{IJ}$ is a topological term related to the Nieh-Yan invariant, thus it will not affect the classical equation of motion in the vacuum case. However, its presence is essential for obtaining the new variables that will be use hereafter.
Varying the action with respect to the connection, $\tfrac{\delta S}{\delta\omega_{a}^{IJ}}=0$, yields the compatibility condition
$P^{KL}{}_{IJ}\mathcal{D}_{b}\left(|e|e_{K}^{[a}e_{L}^{b]}\right)=0\,.$ (4)
This condition expresses the compatibility between the tetrad and the spin connection. This ensures that all tensor fields (such as the Dirac spinor field, the metric, etc.) transform properly under local Lorentz transformations.
To proceed toward a canonical formulation, one foliates the spacetime into spatial slices following the ADM decomposition. For this purpose, a spacetime tensor field, the triad $\mathcal{E}_{I}^{a}=e_{I}^{a}+n_{I}n^{a}$, is introduced where $n^{a}$ is the unit timelike vector field normal to the spatial slices, and $n_{I}\coloneqq e_{I}^{a}n_{a}$. It is also common to fix the boost part of the internal Lorentz transformations by imposing the “time gauge” condition $e_{0}^{a}=n^{a}$ Bojowald:2010qpa .
We can now express the Holst action in terms of the components of the unit normal vector
$n^{a}=N^{-1}(t^{a}-N^{a})$ (5)
where $N$ is the lapse function and $N^{a}$ the shift vector. Making use of the relation $|\text{det}\,e|=N\sqrt{\text{det}\,h}$, where $\text{det}\,h$ is the determinant of the spatial metric $h_{ab}$ induced on the spatial hypersurfaces, the action becomes
$S=\frac{1}{2\kappa}\int d^{4}x\sqrt{\text{det}\,h}P^{IJ}{}_{KL}F_{ab}^{KL}(\omega)(-2t^{a}n_{I}+N\mathcal{E}_{I}^{a}+2N^{a}n_{I})\mathcal{E}_{J}^{b}\,.$ (6)
As we will see below, the first term of the action leads to the symplectic term from which the canonical variables can be read off, plus two constraints (which both turn out to be the same, namely the Gauss constraint). The second and third terms will yield the Hamiltonian and the spatial diffeomorphism constraints, respectively.
Using the definition of the Lie derivative given by $\mathcal{L}_{t}X_{b}^{i}=t^{a}\partial_{a}X_{b}^{i}+X_{a}^{i}\partial_{b}t^{a}$, the first term of the above action can be further decomposed as
$\frac{1}{\kappa\beta}\int d^{4}xE_{j}^{b}\mathcal{L}_{t}\left(\beta K_{b}^{j}+\Gamma_{b}^{j}\right)+\frac{1}{\kappa\beta}\int d^{4}x\Bigl{\{}\beta\omega_{t}^{0j}\left(\partial_{b}E_{j}^{b}+\epsilon_{jk}\,^{l}\Gamma_{b}^{k}E_{l}^{b}-\frac{1}{\beta}\epsilon_{jk}{}^{l}K_{b}^{k}E_{l}^{b}\right)$ (7)
$-\frac{1}{2}\epsilon_{kl}^{j}\omega_{t}^{kl}\left(\partial_{b}E_{j}^{b}+\beta\epsilon_{jm}{}^{n}K_{b}^{m}E_{n}^{b}+\epsilon_{jm}{}^{n}\Gamma_{b}^{m}E_{n}^{b}\right)\Bigr{\}}\,,$
where we have introduce the densitized triad, the extrinsic curvature and the spin connection compatible with the triad, respectively as
$E_{i}^{a}=$ $\sqrt{\text{det}\,h}\,\mathcal{E}_{i}^{a}\,,$ (8)
$K_{a}^{i}=$ $\omega_{a}^{0i}\,,$ (9)
$\Gamma_{a}^{i}=$ $-\frac{1}{2}\epsilon^{i}{}_{jk}\omega_{a}^{jk}\,,$ (10)
with the choice $\epsilon^{0ijk}=\epsilon_{tabc}=1$ for the Levi-Civita symbol. It is now seen that the first term of (7) is the symplectic term of the action from which we can read off the configuration variable, the Ashtekar-Barbero connection, defined as
$A_{a}^{i}=\beta K_{a}^{i}+\Gamma_{a}^{i}\,,$ (11)
and its conjugate momentum the densitized triad $E_{i}^{a}$. Using this definition and with further manipulation, Eq. (7) becomes
$\frac{1}{\kappa\beta}\int d^{4}xE_{j}^{b}\mathcal{L}_{t}A_{b}^{j}+\frac{1}{\kappa\beta}\int d^{4}x\Bigl{\{}\beta\omega_{t}^{0j}\left(\mathcal{D}_{b}^{(A)}E_{j}^{b}-\frac{1+\beta^{2}}{\beta^{2}}\epsilon_{jk}{}^{l}K_{b}^{k}E_{l}^{b}\right)-\frac{1}{2}\epsilon_{kl}^{j}\omega_{t}^{kl}\mathcal{D}_{b}^{(A)}E_{j}^{b}\Bigr{\}}$ (12)
$=\frac{1}{\kappa\beta}\int d^{4}xE_{j}^{b}\mathcal{L}_{t}A_{b}^{j}+\frac{1}{\kappa\beta}\int d^{4}x\Bigl{\{}\Lambda^{j}\mathcal{G}_{j}-(1+\beta^{2})\omega_{t}^{0j}\epsilon_{jk}{}^{l}K_{b}^{k}E_{l}^{b}\Bigr{\}}.$
Here the covariant derivative with respect to the Ashtekar-Barbero connection is defined as $\mathcal{D}_{a}^{(A)}X_{i}^{a}=\partial_{a}X_{i}^{a}+\epsilon_{ij}{}^{k}A_{a}^{j}X_{k}^{a}$ and we have introduced the quantity $\Lambda^{i}=\beta\omega_{t}^{0i}-\frac{1}{2}\epsilon_{jk}^{i}\omega_{t}^{jk}$. Since neither $\Lambda^{i}$ nor $\omega_{t}{}^{0i}$ appear with time derivatives in the action, their conjugate momenta must vanish. Therefore, they serve as Lagrange multipliers of the following constraints
$\mathcal{G}_{i}=$ $\frac{1}{\kappa\beta}\mathcal{D}_{a}^{(A)}E_{i}^{a}\approx 0,$ $\mathcal{S}_{i}=$ $\frac{1}{\kappa\beta}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}\approx 0.$ (13)
The $\approx$” symbol here indicates weak equality [^1] In general, a function $f$ is said to be weakly equal to a function $g$, denoted $f\approx g$, if they agree on the subspace defined by the primary constraints.
To ensure consistency, we must impose the compatibility condition of the densitized triad with the spin connection, which yields the Gauss constraint as
$\mathcal{G}_{i}=\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}\,.$ (14)
As a result, the constraint $\mathcal{S}_{i}$ becomes proportional to $\mathcal{G}_{i}$, indicating that $\mathcal{S}_{i}$ is redundant and can be consistently discarded.
The remaining two terms of the action yeild the Hamiltonian and the spatial diffeomorphism constraints, respectively
$\mathcal{H}=\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\Bigl{\{}\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\Bigr{\}}\,,$ (15)
and
$\mathcal{H}_{a}=\frac{1}{\kappa\beta}E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-\frac{(1+\beta^{2})}{\kappa\beta}\epsilon^{j}{}_{kl}E_{j}^{b}K_{a}^{k}K_{b}^{l}\,,$ (16)
in which
$\mathcal{F}_{ab}^{i}(A)=2\partial_{[a}\Gamma_{b]}^{i}+\epsilon^{i}{}_{jk}\Gamma_{a}^{j}\Gamma_{b}^{k}+2\beta\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{i}+\beta^{2}\epsilon^{i}{}_{jk}K_{a}^{j}K_{b}^{k}$ (17)
is the 2-curvature of the Ashtekar connection and $\mathcal{D}^{(\Gamma)}$ is the covariant derivative with respect to the spin connection. Thus, the full Hamiltonian for a system in which only the gravitational degrees of freedom are present takes the form
$H=\int d^{3}x\left(-\Lambda^{i}\mathcal{G}_{i}+N\mathcal{H}+N^{a}\mathcal{H}_{a}\right)\,.$ (18)
As is well-known, this is the sum of three types of constraints
$\mathcal{G}_{i}=$ $\frac{1}{\kappa\beta}\mathcal{D}_{a}^{(A)}E_{i}^{a}=\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}\,,$ (19)
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\Bigl{\{}\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}\Bigr{\}}\,,$ (20)
$\mathcal{H}_{a}$ $=\frac{1}{\kappa\beta}E_{j}^{b}\mathcal{F}_{ab}^{j}(A)\,.$ (21)
It is worth noting that, on the constraint surface, the Gauss constraint vanishes weakly. This implies that, on this surface, the antisymmetric part of the extrinsic curvature $K_{ab}$ vanishes, rendering $K_{ab}$ symmetric in its indices, as expected.
Proceeding further, additional constraints can be derived by varying the action with respect to suitable combinations of the non-dynamical variables. Notably, this procedure yields an explicit expression for the spin connection $\Gamma_{a}^{i}$ in terms of the triad:
$\Gamma_{a}^{i}=\frac{1}{2}\epsilon^{ijk}e_{k}^{b}\left(\partial_{a}e_{bj}-\partial_{b}e_{aj}+e_{a}^{l}e_{j}^{c}\partial_{c}e_{bl}\right)\,.$ (22)
This equation ensure that the connection $\Gamma_{a}^{i}$ is precisely the spin connection compatible with the triad.
3 Fermions
3.1 Weyl Fermions
We now study how the inclusion of Weyl fermions alters the familiar vacuum scenario. This system has been previously studied in the literature Bojowald:2010qpa ; Das:2008r ; Thiemann:1998QSDV ; Jacobson:1988qta ; Mercuri:2006wb ; PhysRevD.73.084016 ; Date:2011rd . Notably, in Bojowald:2010qpa and Das:2008r , the Dirac matrices themselves were modified to accommodate the curved spacetime geometry. In contrast, the present work adopts an approach in which instead of altering the matrices, we modify the algebraic identities involving them (see Appendix A). This results in a formulation that more closely mirrors the flat spacetime treatment of fermionic fields, allowing a more straightforward comparison between the two cases.
The starting point is the study of the Dirac action discussed in Appendix A.3, specifically Eq.(183). In order to derive the constraints associated to the fermionic contribution, we must revisit the canonical decomposition $\mathcal{E}_{I}^{a}=e_{I}^{a}+n^{a}n_{I}$ where $n_{I}=-\delta_{I}^{0}$ (in the time gauge), as was done in the vacuum case. As a result, the fermionic action becomes
$S_{F}=\frac{i}{2}\int$ $d^{4}x\sqrt{\text{det}\,h}\Bigl{[}\overline{\Psi}\gamma^{I}\left(N\mathcal{E}_{I}^{a}-t^{a}n_{i}+N^{a}n_{I}\right)\mathfrak{D}_{a}\Psi$
$-\overline{\mathfrak{D}_{a}\Psi}\gamma^{I}\left(N\mathcal{E}_{I}^{a}-t^{a}n_{i}+N^{a}n_{I}\right)\Psi\Bigr{]}\,.$ (23)
Once again we encounter three types of terms involving $t^{a},\,N$, and $N^{a}$ The term containing $t^{a}$ leads to
$\frac{i}{2}\int d^{4}x\sqrt{\text{det}\,h}\Bigl{[}\Psi^{\dagger}\partial_{t}\Psi-(\partial_{t}\Psi^{\dagger})\Psi+\omega_{t}^{ij}\Psi^{\dagger}\sigma_{ij}\Psi\Bigr{]}\,,$ (24)
where the first two terms contribute to the symplectic term, while the last one yields the fermionic contribution to the Gauss constraint. As introduced in Section 2, the Gauss Lagrange multiplier is defined as
$\Lambda^{i}=\beta\omega_{t}{}^{0i}-\frac{1}{2}\epsilon^{i}{}_{jk}\omega_{t}{}^{jk}\,.$ (25)
From this, we can obtain $\omega_{t}^{ij}$ as
$\omega_{t}^{ij}=\beta\epsilon^{ij}{}_{k}\omega_{t}{}^{0i}-\epsilon^{ij}{}_{k}\Lambda^{k}\,.$ (26)
Replacing this into (24) allows us to identify the fermionic contributions to both the Gauss constraint and the $\mathcal{S}_{i}$ constraints, respectively
$\mathcal{G}_{i}^{F}=\frac{1}{2}\sqrt{\text{det}\,h}J_{i}\,,$ $\mathcal{S}_{i}^{F}=-\frac{1}{2}\beta\sqrt{\text{det}\,h}J_{i}\,,$ (27)
where the superscript $F$ denotes the fermionic contribution. Here $J^{I}$ is the fermionic axial 4-current defined as $J^{I}=\overline{\Psi}\gamma_{5}\gamma^{I}\Psi$, whose components in terms of left- and right-chiral spinors and Pauli matrices are
$J^{i}=$ $\overline{\Psi}\gamma_{5}\gamma^{i}\Psi=-\Psi_{R}^{\dagger}\sigma^{i}\Psi_{R}-\Psi_{L}^{\dagger}\sigma^{i}\Psi_{L}\,,$ (28)
$J^{0}=$ $\overline{\Psi}\gamma_{5}\gamma^{0}\Psi=-\Psi_{R}^{\dagger}\Psi_{R}+\Psi_{L}^{\dagger}\Psi_{L}\,.$ (29)
The terms in the fermionic action involving the lapse $N$ and the shift vector $N^{a}$ yield, respectively, the fermionic contributions to the Hamiltonian and spatial diffeomorphism constraints as
$\mathcal{H}^{F}=$ $-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{i}\Psi\right)\,,$ (30)
$\mathcal{H}_{a}^{F}=$ $\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{0}\Psi\right)\,.$ (31)
By minimally coupling the vacuum action (see Section 2) to the fermionic action, we obtain the total Hamiltonian of the system, comprising both the vacuum and the (Weyl) fermionic field contributions.
$H_{\text{G+F}}=\int d^{3}x\left(-\Lambda^{i}\mathcal{G}_{i}^{\text{G+F}}-\omega_{t}^{0i}\mathcal{S}_{i}^{\text{G+F}}+N\mathcal{H}^{\text{G+F}}+N^{a}\mathcal{H}_{a}^{\text{G+F}}\right)\,,$ (32)
where
$\mathcal{G}_{i}^{\text{G+F}}=$ $\frac{1}{\kappa\beta}\mathcal{D}_{a}^{(A)}E_{i}^{a}+\frac{1}{2}\sqrt{\text{det}\,h}J_{i},$ (33)
$\mathcal{S}_{i}^{\text{G+F}}=$ $-\frac{1+\beta^{2}}{\kappa\beta}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{1}{2}\beta\sqrt{\text{det}\,h}\,J_{i},$ (34)
$\mathcal{H}^{\text{G+F}}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\Bigl{\{}\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\Bigr{\}}$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{i}\Psi\right),$ (35)
$\mathcal{H}_{a}^{\text{G+F}}=$ $\frac{1}{\kappa\beta}\left(E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon^{j}{}_{kl}E_{j}^{b}K_{a}^{k}K_{b}^{l}\right)-\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}\mathfrak{D}_{a}\Psi+\overline{\mathfrak{D}_{a}\Psi}\gamma^{0}\Psi\right).$ (36)
For the sake of notational simplicity, we will omit the superscript “G+F” in the remaining of this section. Unless otherwise stated, all expressions should be understood as referring to the total system, including both the vacuum and the (Weyl) fermionic fields.
The presence of the fermionic field modifies some of the fundamental relations, in particular, the compatibility condition. This can be seen by varying the combined vacuum and fermionic actions with respect to the connection 1-form $\omega_{a}^{IJ}$, which yields the following modified relation
$\frac{1}{\kappa}\mathcal{D}_{b}(|e|e_{K}^{[a}e_{L}^{b]})P^{KL}{}_{IJ}-\frac{1}{4}|e|e_{M}^{a}\epsilon^{M}{}_{IJN}J^{N}=0,$ (37)
which consequently leads to
$\mathcal{D}_{b}(|e|e_{K}^{[a}e_{L}^{b]})=\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}|e|\left(\epsilon_{KL}{}^{MN}e_{M}^{a}J_{N}-\frac{1}{\beta}\left(e_{K}^{a}J_{L}-e_{L}^{a}J_{K}\right)\right).$ (38)
This implies that the system is no longer torsionless and the canonical variables of the theory now carry an implicit dependence on the fermionic field. Moreover, as already noted in Perez:2005pm , unlike the vacuum case, where the Barbero-Immirzi parameter plays no role in the equations of motion, here the torsion term is proportional to this parameter, effectively governing the strength of the interaction between the gravitational and fermionic fields. As a result, to enforce the compatibility condition in the presence of fermions, the connection 1-form must be modified as
$\omega_{a}^{IJ}=$ $\tilde{\omega}_{a}^{IJ}+C_{a}^{IJ}$ (39)
$=$ $\tilde{\omega}_{a}^{IJ}-\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\left(\epsilon^{IJ}{}_{KL}e_{a}^{K}J^{L}-\frac{2}{\beta}e_{a}^{[I}J^{J]}\right)\,,$
where $C_{a}^{IJ}$ is the contorsion tensor and $\tilde{\omega}_{a}^{IJ}$ denotes the torsion-free (vacuum) connection 1-form.
Now recall that $\Gamma_{a}^{i}=-\epsilon^{i}{}_{jk}\omega_{a}^{jk}$. This relation implies that any modification to the connection 1-form necessarily alters the spin connection $\Gamma_{a}^{i}$ from its vacuum form. To derive the explicit expression for the modified spin connection in terms of the original (vacuum) spin connection and a correction term involving the fermionic field, we consider the variation of the Hamiltonian $H_{\text{G+F}}$ with respect to the spin connection. This variation yields the symmetric component of the modified spin connection. The antisymmetric component is then obtained directly from the compatibility condition (38). Together, these results uniquely determine the full expression for the modified spin connection as
$\Gamma_{a}^{i}=$ $\tilde{\Gamma}_{a}^{i}+C_{a}^{i}$
$=$ $\frac{1}{2}\epsilon^{ijk}e_{k}^{b}\left(\partial_{a}e_{bj}-\partial_{b}e_{aj}+e_{a}^{l}e_{j}^{c}\partial_{c}e_{bl}\right)+\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\left(e_{a}^{i}J^{0}-\frac{1}{\beta}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right),$ (40)
where $\tilde{\Gamma}_{a}^{i}$ is the spin connection in the vacuum case, and $C_{a}^{i}$ is the spatial projection of the contorsion tensor $C_{a}^{IJ}$. Hereafter, all the torsionless (i.e., the vacuum case) variables will be marked with a tilde, while the variables with implicit fermionic contribution will be written without.
Although the modification to the extrinsic curvature can also be analyzed, it is not immediately relevant for the present discussion and will be addressed at a later stage in the paper. For now, we observe that the redefinition of the spin connection leads to a modified Ashtekar-Barbero connection. Its explicit form, incorporating the fermionic contribution, is given by
$A_{a}^{i}$ $=\beta K_{a}^{i}+\Gamma_{a}^{i}$ (41)
$=\beta K_{a}^{i}+\tilde{\Gamma}_{a}^{i}+\frac{\kappa}{4}\frac{\beta^{2}}{1+\gamma^{2}}\left(e_{a}^{i}J^{0}-\frac{1}{\beta}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\,.$
With the modified Ashtekar-Barbero connection at hand, we are now in a position to express the constraints of the theory, explicitly separating the vacuum contributions from those arising due to the fermionic sector.
Taking into account the above considerations, the modified compatibility condition takes the form
$\partial_{a}E_{i}^{a}+\epsilon_{ij}{}^{k}\Gamma_{a}^{j}E_{k}^{a}-\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\epsilon_{ij}{}^{k}\left(e_{a}^{j}J^{0}-\frac{1}{\beta}\epsilon^{j}{}_{mn}e_{a}^{m}J^{n}\right)E_{k}^{a}=0\,.$ (42)
This condition leads to the following Gauss constraint
$\mathcal{G}_{i}=\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{a}^{k}-\frac{1}{2}\frac{\beta^{2}}{1+\beta^{2}}\sqrt{\text{det}\,h}J_{i}\,$ (43)
Once again, we see that the $\mathcal{S}_{i}$ constraint (34) is proportional to the Gauss constraint and can be neglected. Finally, the first class [^2] A constraint is said to be first class if its Poisson bracket with all other constraints vanishes weakly. constraints in the case in which gravity is minimally coupled to a fermionic field are
$\mathcal{G}_{i}=$ $\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{a}^{k}+\frac{1}{2}\frac{\beta^{2}}{1+\beta^{2}}\sqrt{\text{det}\,h}J_{i}\,,$ (44)
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{i}\Psi\right)\,,$ (45)
$\mathcal{H}_{a}=$ $\frac{1}{\beta\kappa}E_{j}^{b}\left(\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon_{kl}^{j}K_{a}^{k}K_{b}^{l}\right)+\frac{i}{2}\left(\overline{\Psi}\gamma^{0}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{0}\Psi\right)\,,$ (46)
where $\mathcal{F}_{ab}^{i}(A)$ is the 2-curvature of the modified Ashtekar connection (41) and $\mathcal{D}_{a}^{(\Gamma)}$ is the covariant derivative with respect to the new spin connection of gravity plus fermions.
Another modification arises in the definition of the fermionic covariant derivatives, which can now be reexpressed in terms of the covariant derivative associated with the modified Ashtekar connection, along with an additional correction term
$\mathfrak{D}_{a}\Psi$ $=\partial_{a}\Psi+\frac{1}{2}\omega_{a}{}^{IJ}\sigma_{IJ}\Psi$ (47)
$=\partial_{a}\Psi+\omega_{a}{}^{0i}\sigma_{0i}\Psi+\frac{1}{2}\omega_{a}{}^{ij}\sigma_{ij}\Psi$
$=\partial_{a}\Psi+K_{a}^{i}\sigma_{0i}\Psi-\frac{1}{2}\epsilon^{ij}{}_{k}\Gamma_{a}^{k}\sigma_{ij}\Psi$
$=D_{a}^{(A)}\Psi+K_{a}^{i}(1+i\beta\gamma_{5})\sigma_{0i}\Psi\,,$
where we have defined the modified covariant derivative
$D_{a}^{(A)}\Psi=\partial_{a}\Psi-iA_{a}^{i}\gamma_{5}\sigma_{0i}\Psi\,.$ (48)
Accordingly, for the conjugate fermionic field, we find
$\overline{\mathfrak{D}_{a}\Psi}=\overline{D_{a}^{(A)}\Psi}+K_{a}^{i}\Psi^{\dagger}(1-i\beta\gamma_{5})\sigma_{0i}\gamma^{0}$ (49)
Inserting these modified expressions into the Hamiltonian and diffeomorphism constraints allows us to compute the new forms of the constraints, respectively
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\right)-\frac{1}{2}\epsilon^{i}{}_{jk}E_{i}^{a}K_{a}^{j}J^{k}+\frac{\beta}{2}E_{i}^{a}K_{a}^{i}J^{0}$ (50)
and
$\mathcal{H}_{a}=$ $\frac{1}{\beta\kappa}E_{j}^{b}\left(\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon^{j}{}_{kl}K_{a}^{k}K_{b}^{l}\right)$ (51)
$+\frac{i}{2}\left(\overline{\Psi}\gamma^{0}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\gamma^{0}\Psi\right)-\frac{\beta}{2}\sqrt{\text{det}\,h}K_{a}^{i}J_{i}\,.$
In the above formulation, the canonical pair for the fermionic system is $(\Psi,\Pi)$ where $\Pi=i\sqrt{\text{det}\,h}\Psi$. However, this leads to a symplectic term in the action in which there is an additional term implying an imaginary correction to the new Ashtekar connection. This will ultimately lead to a modification of the Poisson brackets. This issue can be resolved by taking into account the half-density fermionic fields Bojowald:2010qpa ; Thiemann:1998QSDV ; Thiemann:1997rq . Thus, the new spinors and their conjugate momenta become
$\xi=$ $4]{\text{det}\,h}\Psi,$ $\pi=$ $i\xi^{\dagger}.$ (52)
As a result, we get the correct symplectic term and consequently the following anti-Poisson brackets for the half-density fermionic fields and their momenta
$\{\xi_{A}(x),\pi_{B}(y)\}_{+}=\delta_{AB}\delta(x-y),$ (53)
where $x,\,y$ are points residing on the same spatial hypersurface. Additionally, the fermionic axial 4-current can also be redefined by introducing the densitized axial 4-current
$\overline{J}^{I}=\sqrt{\text{det}\,h}J^{I}.$ (54)
Its components can be expressed in terms of the half-density spinor fields and the generators $\tau^{i}$ of the $su(2)$ Lie algebra, namely:
$\overline{J}^{i}=$ $\overline{\xi}\gamma_{5}\gamma^{i}\xi=-i\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi=-2\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)\,,$ (55)
$\overline{J}^{0}=$ $\overline{\xi}\gamma_{5}\gamma^{0}\xi=-i\pi\gamma^{0}\gamma_{5}\gamma^{0}\xi=i(\pi_{R}\xi_{R}-\pi_{L}\xi_{L})\,.$ (56)
Moreover, one can show that Thiemann:1998QSDV
$D_{a}^{(A)}\Psi=\frac{1}{\sqrt[4]{\text{det}\,h}}D_{a}^{(A)}\xi\,.$ (57)
We are now in a position to verify whether the equations of motion (EOM) derived from this modified theory are consistent with the usual curved spacetime case. To the best of our knowledge, the EOM for this system have not been previously analyzed, and performing this consistency check is essential to ensure the validity of the underlying theory.
To perform this check, our strategy is to show that the EOM coming from our Hamiltonian formulation match those directly coming from the Euler-Lagrange equations. In curved spacetime, where the covariant derivative is the $D_{a}^{(A)}$ defined in Eq. (48), the fermionic EOM for $\Psi$ derived from the Euler-Lagrange equation become
$ie_{0}^{a}\gamma^{0}D_{a}^{(A)}\Psi-\frac{1}{2}\beta e_{0}^{a}K_{a}^{i}\gamma_{5}\gamma_{i}\Psi+ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\Psi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\Psi-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma^{0}\Psi=0\,,$ (58)
which reduces to the well-known EOM for fermionic field in flat spacetime, namely the Dirac equation. Replacing $\Psi$ with its expression in terms of the half-density spinor field $\xi$, and taking into account the spacetime foliation adopted from the beginning, the above EOM becomes
$(t^{a}-N^{a})[i\gamma^{0}D_{a}^{(A)}\xi-\frac{1}{2}\beta K_{a}^{i}\gamma_{5}\gamma_{i}\xi]-N[ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\xi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\xi-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma_{0}\xi]=0\,.$ (59)
On the other hand, the EOM coming from the Hamiltonia formulation can be derived using the anti-Poisson brackets as
$\dot{\xi}=\left\{\xi(x),H_{\text{G+F}}\right\}$ (60)
where the Hamiltonian and diffeomorphism constraints in the above expression written in terms of half-density fields have the following forms
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$ (61)
$-\frac{1}{2}e_{i}^{a}\left(\pi\gamma^{0}\gamma^{i}D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\gamma^{0}\gamma^{i}\xi\right)+\frac{i}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\pi\gamma^{0}\gamma^{k}\xi-\frac{i\beta}{2}e_{i}^{a}K_{a}^{i}\pi\gamma^{0}\gamma_{5}\gamma^{0}\xi\,,$
$\mathcal{H}_{a}=$ $\frac{1}{\kappa\beta}E_{j}^{b}\left(\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon_{kl}^{j}K_{a}^{k}K_{b}^{l}\right)+\frac{1}{2}\left(\pi D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\xi\right)+\frac{i\beta}{2}K_{a}^{i}\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi\,.$
A straightforward computation shows that by keeping the dependence of $A_{a}^{i}$ and $\Gamma_{a}^{i}$ on the fermionic field implicit, these constraints lead to the exact equations of motion as (59).
Finally, we can simplify the constraints by in terms of the new Ashtekar connection, the half-density fermionic fields and the Pauli basis. This yields:
$\mathcal{G}_{i}$ $=\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{\beta^{2}}{1+\beta^{2}}\left(\pi_{L}\tau_{i}\xi_{L}+\pi_{R}\tau_{i}\xi_{R}\right)\,,$ (62)
$\mathcal{H}$ $=\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}\right)+\frac{1+\beta^{2}}{\beta^{2}}\mathcal{D}_{a}^{(\Gamma)}(e_{i}^{a}\mathcal{G}^{i})$ (63)
$+\frac{1}{2}(-1+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}\left(\pi_{L}\tau_{i}\xi_{L}+\pi_{R}\tau_{i}\xi_{R}\right)\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)$
$+\beta e_{i}^{a}D_{a}^{(A)}\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)-\frac{\kappa}{2}(1+2\beta^{2})\mathcal{G}_{i}\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)$
$-ie_{i}^{a}\left(\pi_{R}D_{a}^{(A)}\xi_{R}-\pi_{L}D_{a}^{(A)}\xi_{L}-D_{a}^{(A)}\pi_{R}\xi_{R}+D_{a}^{(A)}\pi_{L}\xi_{L}\right)\,,$ (64)
$\mathcal{H}_{a}$ $=\frac{1}{\kappa\beta}E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-\frac{1+\beta^{2}}{\beta}\mathcal{G}_{i}K_{a}^{i}+\frac{1}{2}\left(\pi_{R}D_{a}^{(A)}\xi_{R}+\pi_{L}D_{a}^{(A)}\xi_{L}-D_{a}^{(A)}\pi_{R}\xi_{R}-D_{a}^{(A)}\pi_{L}\xi_{L}\right)\,.$ (65)
From this form of the constraints, we can easily recognize the vacuum contributions and the added fermionic ones, where now all the variables depend either implicitly or explicitly on the fermionic field. When fermionic fields are absent, the standard vacuum constraints are readily recovered.
3.2 Dirac fermions
Before analyzing the QED framework (i.e., the case involving both fermions and photons and their interaction), it is necessary to study Dirac fermions. All electrically charged fermions are described by Dirac spinors. In contrast, Majorana fermions —being their own antiparticles— must be electrically neutral. Since QED describes interactions between charged particles mediated by photon exchange, the theory inherently involves Dirac fermions. Consequently, we adopt the four-component representation of fermionic fields Pal:2011pb ; Schwartz:2014sze . For further details, refer to Appendix A.4.
The starting point is the following Dirac action
$S_{F}=\int d^{4}x|e|\Bigl{\{}\frac{i}{2}\left(e_{I}^{a}\overline{\Psi}\gamma^{I}\mathfrak{D}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}\gamma^{I}\Psi\right)-m\overline{\Psi}\Psi\Bigr{\}}\,.$ (66)
where the covariant derivatives are defined in Eqs. 181-182. Since the mass term does not contain any derivatives, performing the same computation as in the massless case reveals that the only modification appears in the Hamiltonian constraint, as we will see below. Thus, the constraints for a system in which vacuum is minimally coupled to a massive fermionic field are
$\mathcal{G}_{i}=$ $\frac{1}{\kappa\beta}\mathcal{D}_{a}^{(A)}E_{i}^{a}+\frac{1}{2}\sqrt{\text{det}\,h}J_{i}\,,$ (67)
$\mathcal{S}_{i}=$ $-\frac{1+\beta^{2}}{\kappa\beta}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{1}{2}\beta\sqrt{\text{det}\,h}\,J_{i}\,,$ (68)
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\Bigl{\{}\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\Bigr{\}}$ (69)
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{i}\Psi\right)+\sqrt{\text{det}\,h}\,m\,\overline{\Psi}\Psi\,,$ (70)
$\mathcal{H}_{a}=$ $\frac{1}{\kappa\beta}\left(E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon^{j}{}_{kl}E_{j}^{b}K_{a}^{k}K_{b}^{l}\right)-\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}\mathfrak{D}_{a}\Psi+\overline{\mathfrak{D}_{a}\Psi}\gamma^{0}\Psi\right)\,.$ (71)
In the previous expressions we have once again suppressed the superscript “G+F”.
Following the same steps as in the masseless case, we find that the compatibility condition remains the same as in the Weyl fermions case. As a result, we obtain the same modified spin connection which leads to the same new Ashtekar connection as given in Eq.(41). Hence, writing the fermionic convariant derivatives in terms of the new Ashtekar connection as in Eqs. (47)-(49) leads to the following constraints
$\mathcal{G}_{i}$ $=\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{a}^{k}+\frac{1}{2}\frac{\beta^{2}}{1+\beta^{2}}\sqrt{\text{det}\,h}J_{i}\,,$ (72)
$\mathcal{H}$ $=\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\right)-\frac{1}{2}\epsilon^{i}{}_{jk}E_{i}^{a}K_{a}^{j}J^{k}+\frac{\beta}{2}E_{i}^{a}K_{a}^{i}J^{0}+\sqrt{\text{det}\,h}\,m\,\overline{\Psi}\Psi\,,$ (73)
$\mathcal{H}_{a}$ $=\frac{1}{\beta\kappa}E_{j}^{b}\left(\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon_{kl}^{j}K_{a}^{k}K_{b}^{l}\right)+\frac{i}{2}\left(\overline{\Psi}\gamma^{0}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\gamma^{0}\Psi\right)$
$-\frac{\beta}{2}\sqrt{\text{det}\,h}K_{a}^{i}J_{i}\,,$ (74)
where, as mentioned before, the only difference with respect to the masseless case is the presence of a mass term in the Hamiltonian constraint.
We can once again perform a consistency check by comparing the EOM derived from the Euler-Lagrange equations with the ones derived from the Hamiltonian in the above form. The Euler-Lagrange equations yield gives
$ie_{0}^{a}\gamma^{0}D_{a}^{(A)}\Psi-\frac{1}{2}\beta e_{0}^{a}K_{a}^{i}\gamma_{5}\gamma_{i}\Psi+ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\Psi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\Psi-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma_{0}\Psi-m\Psi=0.$ (75)
By following the same steps as in the massless case, it is straightforward to show that the expression for the fermionic EOM in the Hamiltonian formulation matches this expression.
In contrast to the massless case, it is not convenient to use the bispinor notation and express the constraints in terms of the Pauli basis (see Appendix A.4 for further details on QED and chirality). We will instead present the constraints in terms of the half-density fields as follows
$\mathcal{G}_{i}$ $=\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{i}{2}\frac{\beta^{2}}{1+\beta^{2}}\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi\,,$ (76)
$\mathcal{H}$ $=\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}\right)+\frac{1+\beta^{2}}{\beta^{2}}\mathcal{D}_{a}^{(\Gamma)}(e_{i}^{a}\mathcal{G}^{i})$ (77)
$+\frac{i}{2}\beta e_{i}^{a}D_{a}^{(A)}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-\frac{1}{8}(-1+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}(\pi\gamma^{0}\gamma_{5}\gamma_{i}\xi)(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)$
$-\frac{i\kappa}{4}(1+2\beta^{2})\mathcal{G}_{i}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-\frac{1}{2}e_{i}^{a}\left(\pi\gamma^{0}\gamma^{i}D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\gamma^{0}\gamma^{i}\xi\right)-im\pi\gamma^{0}\,,$ (78)
$\mathcal{H}_{a}$ $=\frac{1}{\kappa\beta}E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-\frac{1+\beta^{2}}{\beta}\mathcal{G}_{i}K_{a}^{i}+\frac{1}{2}\left(\pi D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\xi\right)\,.$ (79)
4 QED
4.1 The full Hamiltonian
In this section, we examine a system in which photons —denoted by a $\gamma$ with no indices— are coupled to a fermionic field, and both fields are minimally coupled to gravity. Unlike the standard formulation of QED where the background spacetime is fixed, here the background is allowed to be dynamical. To maintain clarity and avoid an overload of calculations at once, we begin by analyzing the QED action in curved spacetime using modified Ashtekar-Barbero variables, an approach which, to the best of our knowledge, has not been previously explored in the literature. After deriving the QED contributions to the constraints, we will then proceed to minimally couple the full system to the gravitational field.
The QED action in curved spacetime takes the form:
$S_{\text{QED}}=$ $S_{F}+S_{\gamma}$ (80)
$=$ $\int d^{4}x|e|\Bigl{\{}\frac{i}{2}\left(e_{I}^{a}\overline{\Psi}\gamma^{I}\tilde{\mathfrak{D}}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}\gamma^{I}\Psi\right)-m\overline{\Psi}\Psi\Bigr{\}}$
$+\int d^{4}\sqrt{-\text{det}\,g}\left(-\frac{1}{4}g^{ac}g^{bd}F_{cd}F_{ab}\right)\,.$
where the fermionic covariant derivatives contain an interaction term between the photon and the fermionic field, as follows
$\tilde{\mathfrak{D}}_{a}\Psi=$ $\partial_{a}\Psi+\frac{1}{2}\omega_{a}{}^{IJ}\sigma_{IJ}\Psi+i\,q\,A_{a}\Psi\,,$ (81)
$\overline{\tilde{\mathfrak{D}}_{a}\Psi}=$ $\partial_{a}\overline{\Psi}+\frac{1}{2}\omega_{a}{}^{IJ}\Psi^{\dagger}\sigma_{IJ}^{\dagger}\gamma^{0}-i\,q\,A_{a}\overline{\Psi}\,.$ (82)
In Eq. (80) and in the following, $A_{a}$ denotes the electromagnetic 4-vector potential, $q$ the electric charge and $F_{ab}=(\nabla_{a}A_{b}-\nabla_{b}A_{a})$ is the electromagnetic field tensor with $\nabla_{a}$ being the covariant derivative with respect to the Christoffel symbol.
Let us begin by focusing solely on the photonic contribution, described by the action $S_{\gamma}$. Given that the spacetime metric decomposes as $g^{ab}=h^{ab}-n^{a}n^{b}$, its contraction with the EM field tensors becomes
$g^{ac}g^{bd}F_{cd}F_{ab}=$ $h^{ac}h^{bd}F_{cd}F_{ab}-\frac{2}{N^{2}}h^{bd}(t^{a}-N^{a})(\nabla_{a}A_{b}-\nabla_{b}A_{a})(t^{c}-N^{c})$ (83)
×$\times(\nabla_{c}A_{d}-\nabla_{d}A_{c})$
$=$ $h^{ac}h^{bd}F_{cd}F_{ab}-\frac{2}{N^{2}}h^{bd}\left(\dot{A}_{b}-\nabla_{b}(t^{a}A_{a})-N^{a}F_{ab}\right)$
×$\times\left(\dot{A}_{d}-\nabla_{d}(t^{c}A_{c})-N^{c}F_{cd}\right)\,.$
Substituting this expression into the photonic action yields:
$\begin{split}S_{\gamma}=&-\frac{1}{4}\int d^{4}x\sqrt{\text{det}\,h}\Bigl{\{}Nh^{ac}h^{bd}F_{cd}F_{ab}-\frac{2}{N}h^{bd}\left(\dot{A}_{b}-\nabla_{b}(t^{a}A_{a})-N^{a}F_{ab}\right)\\ &\times\left(\dot{A}_{d}-\nabla_{d}(t^{c}A_{c})-N^{c}F_{cd}\right)\Bigr{\}}\,.\end{split}$ (84)
Next, we express this action in the Hamiltonian form as $S_{\gamma}=\int d^{4}x\left(\pi^{a}\dot{A}_{a}-H_{\gamma}\right)$, in which $\pi^{a}$ denotes the momentum conjugate to the spatial components of the gauge field $A_{a}$. This momentum is computed via functional differentiation
$\pi^{a}$ $=\frac{\delta S_{\gamma}}{\delta\dot{A}_{a}}$ (85)
$=\frac{\sqrt{\text{det}\,h}}{N}h^{ab}\left(\dot{A}_{b}-\nabla_{b}(t^{c}A_{c})-N^{c}F_{cb}\right)$
$=\sqrt{\text{det}\,h}h^{ab}n^{c}F_{cb}\,,$
where in the last last we have used part of the second term of (83). From this result, we observe that $\pi^{a}$ is a densitized vector field. Substituting this expression back into the action, we obtain the action in the desired form as
$S_{\gamma}=$ $\int d^{4}x\Bigl{\{}-\sqrt{\text{det}\,h}\frac{N}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{2}\frac{N}{\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}\Bigr{\}}$ (86)
$=$ $\int d^{4}x\bigg{\{}\pi^{a}\dot{A}_{a}+t^{b}A_{b}\nabla_{a}\pi^{a}-N\left(\frac{\sqrt{\text{det}\,h}}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{2\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}\right)$
$-N^{b}F_{ba}\pi^{a}\bigg{\}}\,.$
Once again one finds four terms corresponding to symplectic term, and Gauss, Hamiltonian and spatial diffeomorphism constraints. In particular, since the temporal component of the gauge field $A_{a}$ does not possess a conjugate momentum, it acts as a Lagrange multiplier. Specifically, $t^{b}A_{b}=A_{t}$ enforces a constraint, which leads to the presence of the Gauss constraint associated with $U(1)$ gauge symmetry. The full set of constraints for the electromagnetic field in the Hamiltonian formulation becomes
$\mathcal{G}^{\gamma}$ $=\nabla_{a}\pi^{a}\,,$ (87)
$\mathcal{H}^{\gamma}$ $=\frac{\sqrt{\text{det}\,h}}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{2\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}\,,$ (88)
$\mathcal{H}_{a}^{\gamma}$ $=F_{ab}\pi^{b}\,.$ (89)
Let us now perform the same analysis for the full QED action $S_{\text{QED}}$ in (80). Beginning with the contributions involving the time vector $t^{a}$, we identify the following terms
$\begin{split}\int d^{4}x&\Bigl{[}\pi^{a}\dot{A}_{a}+\frac{i}{2}\sqrt{\text{det}\,h}\left(\Psi^{\dagger}\partial_{t}\Psi-(\partial_{t}\Psi)\Psi\right)\Bigr{]}\\ &+\int d^{4}x\left\{\frac{i}{2}\sqrt{\text{det}\,h}\,\omega_{t}{}^{ij}\Psi^{\dagger}\sigma_{ij}\Psi+A_{t}\left(\nabla_{a}\pi^{a}-\sqrt{\text{det}\,h}\,q\Psi^{\dagger}\Psi\right)\right\}\,.\end{split}$ (90)
We observe that the first integral corresponds to the symplectic term, while the second integral contains two Gauss constraints, each multiplied by a Lagrange multiplier, respectively $\omega_{t}^{ij}$ and $A_{t}$. One of these constraints depends solely on internal indices and represents the fermionic contribution to the vacuum Gauss constraint. The other involves only spacetime indices —specifically, the spatial ones— and reduces to Gauss’ Law in the absence of the fermionic field. Hence, the full QED contribution to the Gauss constraint is
$\Lambda\mathcal{G}^{\text{QED}}=\omega_{t}{}^{ij}\,\frac{i}{2}\sqrt{\text{det}\,h}\,\Psi^{\dagger}\sigma_{ij}\Psi+A_{t}\left(\nabla_{a}\pi^{a}-\sqrt{\text{det}\,h}\,q\Psi^{\dagger}\Psi\right)\,.$ (91)
As in the previous section, the Lagrange multiplier $\omega_{t}^{ij}$ will be expressed in terms of $\Lambda_{i}$ and $\omega_{t}^{0i}$ in order to derive its contribution to the vacuum Gauss and $\mathcal{S}_{i}$ constraints.
Next, considering the terms involving the lapse function $N$, we obtain the QED Hamiltonian constraint as
$\begin{split}\mathcal{H}^{\text{QED}}&=\frac{\sqrt{\text{det}\,h}}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{2\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\tilde{\mathfrak{D}}_{a}\Psi-\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{i}\Psi\right)\\ &+\sqrt{\text{det}\,h}\,m\overline{\Psi}\Psi\,.\end{split}$ (92)
Finally, the terms multiplied by the shift vector $N^{a}$ yield the QED diffeomorphism constraint
$\mathcal{H}_{a}^{\text{QED}}=F_{ab}\pi^{b}+\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}\tilde{\mathfrak{D}}_{a}\Psi-\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{0}\Psi\right)\,.$ (93)
It is now possible to minimally couple the gravitational field to the photonic and fermionic fields in order to obtain the full constraints of the system, which are
$\Lambda^{i}\mathcal{G}_{i}^{\text{G+F}}+A_{t}\mathcal{G}^{\text{QED}}=$ $\left(\beta\omega_{t}{}^{0i}-\frac{1}{2}\epsilon^{i}{}_{jk}\omega_{t}{}^{jk}\right)\Bigl{[}\frac{1}{\kappa\beta}\left(\partial_{a}E_{i}^{a}+\epsilon_{ip}{}^{q}A_{a}^{p}E_{q}^{a}\right)+\frac{1}{2}\sqrt{\text{det}\,h}J_{i}\Bigr{]}$
$+A_{t}\left(\nabla_{a}\pi^{a}-\sqrt{\text{det}\,h}q\Psi^{\dagger}\Psi\right)\,,$ (94)
$\mathcal{S}_{i}^{\text{Full}}=$ $-\frac{1+\beta^{2}}{\kappa\beta}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{1}{2}\beta\sqrt{\text{det}\,h}J_{i}\,,$
$\mathcal{H}^{\text{Full}}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
$+\frac{\sqrt{\text{det}\,h}}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{2\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\tilde{\mathfrak{D}}_{a}\Psi-\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{i}\Psi\right)+\sqrt{\text{det}\,h}m\overline{\Psi}\Psi\,,$ (95)
$\mathcal{H}_{a}^{\text{Full}}=$ $\frac{1}{\kappa\beta}E_{i}^{b}\left(\mathcal{F}_{ab}^{i}(A)-(1+\beta^{2})\epsilon^{i}{}_{jk}K_{a}^{j}K_{b}^{k}\right)$
$+\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}\tilde{\mathfrak{D}}_{a}\Psi-\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{0}\Psi\right)+F_{ab}\pi^{b}\,.$ (96)
None of the terms involving the 4-vector potential $A_{a}$ couples to the Lorentz connection 1-form $\omega_{a}^{IJ}$. As a result, varying the full action, where the gravitational field is coupled to both fermionic and photonic fields, with respect to the 1-form yields the same result as in the case where only gravity and the fermionic field are present,
$\frac{\delta S_{\text{Full}}}{\delta\omega_{a}{}^{IJ}}=\frac{\delta S_{F}}{\delta\omega_{a}{}^{IJ}}=-\frac{1}{4}|e|e_{K}^{a}\epsilon^{K}{}_{IJL}J^{L}\,.$ (97)
Thus, as expected, photons do not imply torsion. Hence, the compatibility condition remains the following
$\mathcal{D}_{a}\left(|e|e^{[c}{}_{K}e^{a]}{}_{L}\right)=\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}|e|\Bigl{[}\epsilon_{KL}{}^{MN}e_{M}^{c}J_{N}-\frac{1}{\beta}\left(e_{K}^{c}J_{L}-e_{L}^{C}J_{K}\right)\Bigr{]}\,.$ (98)
Consequently, the new spin connection takes the same form as in Eq.(40). Moreover, the full constraints in terms of the modified Ashtekar connection look like Eqs. (94)-(96), but now $A_{a}^{i}$ implicitly depends on the fermionic field given in Eq. (41).
Following the steps we performed before, we can now reformulate the fermionic covariant derivatives in terms of the modified Ashtekar connection, making its dependence on this connection explicit. They are written as follows
$\tilde{\mathfrak{D}}_{a}\Psi=$ $D_{a}^{(A)}\Psi+K_{a}^{i}\left(1+i\beta\gamma_{5}\right)\sigma_{0i}\Psi+iqA_{a}\Psi\,,$ (99)
$\overline{\tilde{\mathfrak{D}}_{a}\Psi}=$ $\overline{D_{a}^{(A)}\Psi}+K_{a}^{i}\Psi^{\dagger}\left(1-i\beta\gamma_{5}\right)\sigma_{0i}\gamma^{0}-iqA_{a}\overline{\Psi}\,,$ (100)
where $D_{a}^{(A)}$ has been defined in Eq. (48). Hence, the Hamiltonian constraint become
$\mathcal{H}^{\text{Full}}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
$+\frac{\sqrt{\text{det}\,h}}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{2\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\gamma^{i}\Psi\right)$
$-\frac{1}{2}\epsilon^{i}{}_{jk}E_{i}^{a}K_{a}^{j}J^{k}+\frac{\beta}{2}E_{i}^{a}K_{a}^{i}J^{0}+qE_{i}^{a}A_{a}\mathcal{J}^{i}+\sqrt{\text{det}\,h}\,m\overline{\Psi}\Psi\,,$ (101)
where $\mathcal{J}^{I}=\overline{\Psi}\gamma^{I}\Psi$ is the fermionic 4-current, and the spatial diffeomorphism constraint turns out to be
$\mathcal{H}_{a}^{\text{Full}}=$ $\frac{1}{\kappa\beta}E_{i}^{b}\left(\mathcal{F}_{ab}^{i}(A)-(1+\beta^{2})\epsilon^{i}{}_{jk}K_{a}^{j}K_{b}^{k}\right)+\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\gamma^{0}\Psi\right)$
$-\frac{\beta}{2}\sqrt{\text{det}\,h}K_{a}^{i}J_{i}-\sqrt{\text{det}\,h}qA_{a}\mathcal{J}^{0}+F_{ab}\pi^{b}\,.$ (102)
Next, we turn our attention to the EOM. First, we derive the EOM using the Euler-Lagrange equation. The fermionic covariant derivative remains the same as in the Dirac fermions section. However, an additional term now appears, which is the electromagnetic coupling. Consequently, in curved spacetime, the EOM for the fermionic field $\Psi$ takes the form
$\begin{split}ie_{0}^{a}\gamma^{0}D_{a}^{(A)}\Psi&-\frac{1}{2}\beta e_{0}^{a}K_{a}^{i}\gamma_{5}\gamma_{i}\Psi+ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\Psi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\Psi\\ &-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma^{0}\Psi-m\Psi-qA_{a}e_{0}^{a}\gamma^{0}\Psi-qA_{a}e_{i}^{a}\gamma^{i}\Psi=0\,,\end{split}$ (103)
which is consistent with the well-known QED EOM for fermions. On the other hand, the EOM for the photon field can be obtained by extremizing the full action $S_{\text{Full}}$ with respect to the electromagnetic 4-vector potential $A_{a}$. This yields a background-independent form of Maxwell’s equations. More concretely, the photonic Euler-Lagrange equation reads
$\frac{\delta\mathcal{L}^{\text{QED}}}{\delta A_{b}}-\nabla_{c}\frac{\delta\mathcal{L}^{\text{QED}}}{\delta\left(\nabla_{c}A_{b}\right)}=0\,,$ (104)
which leads to the following expression for the photonic EOM in curved spacetime
$\nabla_{c}\left(g^{ac}g^{bd}F_{ad}\right)=qe_{I}^{b}\mathcal{J}^{I}\,,$ (105)
which is consistent with standard QED theory.
We can rewrite the fermionic EOM in terms of the half-density fermionic field and the foliation of the spacetime introduced from the beginning, obtaining
$(t^{a}-N^{a})$ $\Bigl{[}i\gamma^{0}D_{a}^{(A)}\xi-\frac{\beta}{2}K_{a}^{i}\gamma_{5}\gamma_{i}\xi-qA_{a}\gamma^{0}\xi\Bigr{]}-N\Bigl{[}ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\xi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\xi$
$-\frac{\beta}{2}e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma^{0}\xi-m\xi-qA_{a}e_{i}^{a}\gamma^{i}\xi\Bigr{]}=0\,.$ (106)
On the other hand, in the Hamiltonian formulation, we can substitute the half-density field into the Hamiltonian and diffeomorphism constraints, and then use the fermionic anti-Poisson brackets, to derive the Hamiltonian EOM. Doing so, we recover the same result as obtained through the Euler-Lagrange method in (106).
As for the photonic field, additional manipulation of the equation derived from the Euler-Lagrange method is required before it can be directly compared to the result obtained from the Hamiltonian formalism. By substituting $g^{ab}=h^{ab}-n^{a}n^{b}$ and expressing $\pi^{a}$ in terms of the electromagnetic field tensor $F_{ab}$ (Eq. (85)) in the Euler-Lagrange EOM, the photonic field EOM becomes
$\nabla_{c}\left(h^{ac}h^{bd}F_{ad}\right)+\nabla_{c}\left(\frac{1}{\sqrt{\text{det}\,h}}n^{b}\pi^{c}\right)-\nabla_{c}\left(\frac{1}{\sqrt{\text{det}\,h}}n^{c}\pi^{b}\right)-qe_{I}^{b}\mathcal{J}^{I}=0\,.$ (107)
This is the EOM for the conjugate momentum $\pi^{a}$ of the 4-vector potential $A_{a}$.
On the other hand, in order to obtain the EOM via the Hamiltonian formalism, we need to consider the following photonic Poisson bracket
$\{A_{a}(x),\pi^{b}(y)\}=\delta_{a}^{b}\delta(x-y)\,,$ (108)
which implies
$\dot{\pi}^{a}=-\frac{\delta H^{\text{Full}}}{\delta A_{a}}\,.$ (109)
Substituting the full Hamiltonian in the previous equation, we obtain the expected result, consistent with the Euler-Lagrange method.
Finally, once again we impose the compatibility condition and observe that the $\mathcal{S}_{i}$ constraint is proportional to the Gauss constraint arising from the fermionic modification of the vacuum. As a result, $\mathcal{S}_{i}$ can be discarded. Therefore, the remaining (first class) constraints are the Gauss constraint derived in the previous section and the one associated with the QED sector, together with the Hamiltonian and the diffeomorphism constraints. Moreover, since we are working with Dirac spinors —and more generally within the framework of QED— the considerations regarding chirality remain applicable. Consequently, the final full constraints, expressed in terms of the modified Ashtekar-Barbero connection and the half-density fermionic field, are
$\Lambda^{i}\mathcal{G}_{i}^{\text{G+F}}+A_{t}\mathcal{G}^{\text{QED}}=$ $\Lambda^{i}\Bigl{[}\frac{1}{\kappa}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{i}{2}\frac{\beta^{2}}{1+\beta^{2}}\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi\Bigr{]}+A_{t}\left(\nabla_{a}\pi^{a}+iq\pi\xi\right)\,,$ (110)
$\mathcal{H}^{\text{Full}}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}\right)$
$+\frac{1+\beta^{2}}{\beta^{2}}\mathcal{D}_{a}^{(\Gamma)}\left(e_{i}^{a}\mathcal{G}^{i}\right)+\frac{i}{2}\beta e_{i}^{a}D_{a}^{(A)}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)$
$-\frac{1}{8}(-1+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}(\pi\gamma^{0}\gamma_{5}\gamma_{i}\xi)(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)$
$-\frac{i\kappa}{4}(1+2\beta^{2})\mathcal{G}_{i}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-im\pi\gamma^{0}$
$-\frac{1}{2}e_{i}^{a}\left(\pi\gamma^{0}\gamma^{i}D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\gamma^{0}\gamma^{i}\xi\right)-iqe_{i}^{a}A_{a}\pi\gamma^{0}\gamma^{i}\xi$
$+\frac{\sqrt{\text{det}\,h}}{4}h^{ac}h^{bd}F_{cd}F_{ab}+\frac{1}{\sqrt{\text{det}\,h}}h_{ab}\pi^{a}\pi^{b}\,,$ (111)
$\mathcal{H}_{a}^{\text{Full}}=$ $\frac{1}{\kappa\beta}E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-\frac{1+\beta^{2}}{\beta}\mathcal{G}_{i}K_{a}^{i}+\frac{1}{2}\left(\pi D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\xi\right)$
$+iqA_{a}\pi\xi+F_{ab}\pi^{b}\,.$ (112)
4.2 Parity transformation
In this final subsection, we analyze the behavior of the full system under parity transformation. A similar study was conducted in Das:2008r for a system where gravity is coupled solely to a fermionic field. Here, we extend the analysis to the full system, where the gravitational field is minimally coupled to both fermionic and photonic fields, providing clearer, more detailed, and more readable final results.
To determine the transformation properties of the new Ashtekar connection under parity, we first need to identify the torsion contribution to the extrinsic curvature. We decompose the extrinsic curvature $K_{a}^{i}$ into a torsionless part $\tilde{K}_{a}^{i}$ and a torsion contribution $k_{a}^{i}$, as follows
$K_{a}^{i}=\tilde{K}_{a}^{i}+k_{a}^{i}\,.$ (113)
In order to provide and explicit expression for $k_{a}^{i}$, the starting point is the EOM of the canonical variables $(A_{a}^{i},E_{i}^{a})$, which are
$\mathcal{L}_{t}A_{a}^{i}=$$\left\{A_{a}^{i},H^{\text{\text{Full}}}\right\}=\kappa\beta\frac{\delta H^{\text{\text{Full}}}}{\delta E_{i}^{a}}\,,$ (114)
$\mathcal{L}_{t}E_{i}^{a}=$$\left\{E_{i}^{a},H^{\text{\text{Full}}}\right\}=-\kappa\beta\frac{\delta H^{\text{\text{Full}}}}{\delta A_{a}^{i}}\,,$ (115)
[…]
$\mathcal{L}_{t}E_{p}^{c}=$ $\left(\frac{1}{2}\epsilon^{i}{}_{jk}\omega_{t}^{jk}-\beta\omega_{t}^{0i}\right)\epsilon_{ip}{}^{n}E_{n}^{c}$
$+N\left[\frac{1}{\beta\sqrt{\text{det}\,h}}\left(E_{p}^{c}E_{m}^{a}A_{a}^{m}-E_{m}^{c}E_{p}^{a}A_{a}^{m}\right)-\frac{1}{\beta}\partial_{a}\left(\epsilon^{acb}e_{bp}\right)-\frac{1}{\kappa}\epsilon^{i}{}_{kp}E_{i}^{c}J^{k}\right]$
$+N^{a}\partial_{a}E_{p}^{c}+E_{p}^{c}\partial_{a}N^{a}-E_{p}^{b}\partial_{b}N^{c}-\epsilon^{j}{}_{kp}N^{a}E_{j}^{c}A_{a}^{k}+\frac{\kappa}{\beta}N^{c}\mathcal{G}_{p}^{\text{G+F}}$ (116)
It is worth noting that, as in the vacuum case, the parity of the extrinsic curvature —and, consequently, that of the entire system— can be defined only on the constraint hypersurface. In particular, we must impose the vanishing of the Gauss constraint $\mathcal{G}^{\text{G+F}}$, which will allow us to express the extrinsic curvature in terms of the fermionic axial four-current. However, we will later observe that the same dependence of the extrinsic curvature on the fermionic field could also be derived directly from the modified connection 1-form obtained via the compatibility condition. Nevertheless, this alternative route does not resolve the issue of determining the behavior of the system under parity transformations. Even in that case, one must still restrict to the hypersurface defined by $\mathcal{G}^{\text{G+F}}=0$ in order to properly analyze the parity properties of the variables of the system. Hence, we neglect the last term in the previous expression. Contracting the previous equation with $e_{cq}$ and $\delta^{pq}$, and then symmetrizing the result, we obtain
$\begin{split}e_{cq}\mathcal{L}_{t}E_{p}^{c}&+e_{cp}\mathcal{L}_{t}E_{q}^{c}+N\frac{1}{\beta}\epsilon^{acb}\left(e_{cq}\partial_{a}e_{bp}+e_{cp}\partial_{a}e_{bq}-\delta_{pq}e_{c}^{i}\partial_{a}e_{bi}\right)\\ &+\frac{N}{\beta}\sqrt{\text{det}\,h}\left(e_{p}^{c}A_{cq}+e_{q}^{c}A_{cp}\right)+\sqrt{\text{det}\,h}\partial_{b}N^{c}\left(e_{cq}e_{p}^{b}+e_{cp}e_{q}^{b}\right)=0\,.\end{split}$ (117)
To isolate the extrinsic curvature, we first expand the term involving the Ashtekar connection as
$\begin{split}\left(e_{p}^{c}A_{cq}+e_{q}^{c}A_{cp}\right)=&\beta\left(e_{p}^{c}\tilde{K}_{cq}+e_{q}^{c}\tilde{K}_{cp}\right)+\beta\left(e_{p}^{c}k_{cq}+e_{q}^{c}k_{cp}\right)+\left(e_{p}^{c}\tilde{\Gamma}_{cq}+e_{q}^{c}\tilde{\Gamma}_{cp}\right)\\ &+\left(e_{p}^{c}C_{cq}+e_{q}^{c}C_{cp}\right)\,.\end{split}$ (118)
Then, substituting this expression into Eq. (117) and using the definition of the torsionless spin connection in terms of the triad (22) and the torsion contribution to the spin connection, we find
$\begin{split}e_{cq}\mathcal{L}_{t}E_{p}^{c}&+e_{cp}\mathcal{L}_{t}E_{q}^{c}+\sqrt{\text{det}\,h}\partial_{b}N^{c}\left(e_{cq}e_{p}^{b}+e_{cp}e_{q}^{b}\right)\\ &+\frac{N}{\beta}\sqrt{\text{det}\,h}\left[\beta\left(e_{p}^{c}\tilde{K}_{cq}+e_{q}^{c}\tilde{K}_{cp}\right)+\beta\left(e_{p}^{c}k_{cq}+e_{q}^{c}k_{cp}\right)+\frac{\kappa}{2}\frac{\beta^{2}}{1+\beta^{2}}\delta_{pq}J^{0}\right]=0\,.\end{split}$ (119)
To isolate the torsionless part of the extrinsic curvature, we use its definition
$\tilde{K}_{ab}=\frac{1}{2N}\left(\dot{h}_{ab}-2\tilde{D}_{(a}N_{b)}\right)\,,$ (120)
where $\tilde{D}_{a}N_{b}=h_{a}^{c}h_{b}^{d}\nabla_{c}N_{d}$ with $\nabla_{a}$ the covariant derivative with respect to the Christoffel symbols. Contracting this expression with the triad and the densitized triad and symmetrizing yields
$E_{p}^{c}\tilde{K}_{cq}+E_{q}^{c}\tilde{K}_{cp}=-\frac{1}{N}\left(e_{cq}\mathcal{L}_{t}E_{p}^{c}+e_{cp}\mathcal{L}_{t}E_{q}^{c}+(E_{p}^{c}e_{q}^{d}+E_{q}^{c}e_{p}^{d})\partial_{c}N_{d}\right)\,.$ (121)
Substituting this back into Eq.(119), the symmetric torsion contribution to the extrinsic curvature is found to obey
$e_{p}^{c}k_{cq}+e_{q}^{c}k_{cp}=-\frac{1}{2}\frac{\kappa\beta}{1+\beta^{2}}\delta_{pq}J^{0}\,.$ (122)
For the antisymmetric expression, we can take advantage of the vanishing Gauss constraint $\mathcal{G}^{\text{G+F}}$ and separate the torsionless and torsion contribution to the extrinsic curvature as
$\mathcal{G}_{i}^{\text{G+F}}=\frac{1}{\kappa}\epsilon_{ij}{}^{k}(\tilde{K}_{a}^{j}+k_{a}^{j})E_{k}^{a}+\frac{1}{2}\frac{\beta^{2}}{1+\beta^{2}}\sqrt{\text{det}\,h}J_{i}=0\,,$ (123)
Since we are considering the constraint hypersurface, the torsionless extrinsic curvature is symmetric. Hence, the antisymmetric part of the torsion contribution to the extrinsic obeys
$e_{q}^{c}k_{cp}-e_{p}^{c}k_{cq}=-\frac{1}{2}\frac{\kappa\beta^{2}}{1+\beta^{2}}\epsilon^{i}{}_{pq}J_{i}\,.$ (124)
By summing the symmetric and antisymmetric components, we derive the expression for the torsion contribution to the extrinsic curvature in terms of the fermionic axial 4-current as
$k_{a}^{i}=-\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\left(\frac{1}{\beta}e_{a}^{i}J^{0}+\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\,.$ (125)
As a consequence, the complete splitting of the Ashtekar connection into torsionless and torsion contributions becomes
$\begin{split}A_{a}^{i}&=\left(\beta\tilde{K}_{a}^{i}+\tilde{\Gamma}_{a}^{i}\right)+\left(\beta k_{a}^{i}+C_{a}^{i}\right)\\ &=\left(\beta\tilde{K}_{a}^{i}+\tilde{\Gamma}_{a}^{i}\right)-\frac{\kappa\beta}{4}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\,.\end{split}$ (126)
where $C_{a}^{i}$ denotes the spatial projection of the contorsion tensor as defined in (40). In the second line, we have used the explicit forms of $C_{a}^{i}$ and $k_{a}^{i}$ to express their contribution in terms of the fermionic current.
We can now examine the behavior of the various variables under a parity transformation. It is important to specify whether we are referring to internal or spacetime indices, as the Gauss constraint can be decomposed according to the two distinct Lagrange multipliers, one of which involves only spacetime indices. Under a parity transformation, the components associated with the internal indices of the triad change orientation while the fermionic field transform as
$\mathds{P}e_{i}^{a}=$ $-e_{i}^{a}\,,$ (127)
$\mathds{P}\Psi=$ $\gamma^{0}\Psi\,,$ (128)
where $\mathds{P}$ is the parity operator. Moreover, the components of the fermionic axial 4-current and of the fermionic 4-current transform respectively as
$\mathds{P}J^{0}=$ $-J^{0}\,,$ $\mathds{P}J^{i}=$ $J^{i}\,,$ (129)
and
$\mathds{P}\mathcal{J}^{0}=$ $\mathcal{J}^{0}\,,$ $\mathds{P}\mathcal{J}^{i}=$ $-\mathcal{J}^{i}\,.$ (130)
The torsionless part of the spin connection is proportional to an even number of triads, while the torsion contribution is proportional to the triad and different components of the fermionic axial 4-current. Thus, the spin connection transforms under parity as
$\mathds{P}\Gamma_{a}^{i}=\tilde{\Gamma}_{a}^{i}+\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\left(e_{a}^{i}J^{0}+\frac{1}{\beta}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\,.$ (131)
On the other hand, the torsionless extrinsic curvature is contracted with the triad $\tilde{K}_{a}^{i}=e^{bi}\tilde{K}_{ab}$, hence the extrinsic curvature transform as
$\mathds{P}K_{a}^{i}=-\tilde{K}_{a}^{i}-\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\left(\frac{1}{\beta}e_{a}^{i}J^{0}-\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\,.$ (132)
By combining the transformation rules for triad, spin connection, extrinsic curvature, and fermionic axial 4-current, we derive the transformation rule for the Ashtekar-Barbero connection which becomes
$\mathds{P}A_{a}^{i}=\tilde{\Gamma}_{a}^{i}-\beta\left(\tilde{K}_{a}^{i}-\frac{\kappa}{4}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\,.$ (133)
For transformation of spacetime indices, we must analyze how the covariant derivative involving Christoffel symbols behaves under a parity transformation. To this end, we consider a generic spacetime 4-vector $V^{\mu}$ and apply the parity transformation operator to it as
$\mathds{P}V^{\mu}=P_{\nu}^{\mu}V^{\nu},$ (134)
where $P_{\nu}^{\mu}$ are the components of the parity operator, given by
$P_{\nu}^{\mu}=\begin{pmatrix}1&0&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\end{pmatrix}\,.$ (135)
The transfromation of a covector $V_{\mu}$ under parity is then expressed as
$\mathds{P}V_{\mu}=(\left(P^{-1}\right)_{\mu}^{\nu}V_{\nu}.$ (136)
Accordingly, the covariant derivative involving the Christoffel connection, when applied to a 4-covector, transforms under parity as
$\mathds{P}\left(\nabla_{\mu}V_{\nu}\right)=$ $\left(P^{-1}\right)_{\mu}^{\sigma}\partial_{\sigma}\left[\left(P^{-1}\right)_{\nu}^{\rho}V_{\rho}\right]$
$-\frac{1}{2}g^{\rho\sigma}\left[\left(P^{-1}\right)_{\mu}^{\tau}\partial_{\tau}g_{\sigma\nu}+\left(P^{-1}\right)_{\nu}^{\tau}\partial_{\tau}g_{\sigma\mu}-\left(P^{-1}\right)_{\sigma}^{\tau}\partial_{\tau}g_{\mu\nu}\right]\,.$ (137)
Consequently, if $\mu$ is a spatial index, the transformation rule is $\mathds{P}(\nabla_{a}V_{\nu})=-\nabla_{a}(\mathds{P}V_{\nu})$. Note that we use $a,\,b,\ldots$ indices as abstract indices (either spacetime of spatial) but here we meant a spatial abstract index.
A 4-vector with purely spacetime indices, such as the electromagnetic 4-vector potential $A^{a}$ , transforms under parity as
$\mathds{P}A^{t}=$ $A^{t}\,,$ $\mathds{P}A^{a}=$ $-A^{a}\,,$ (138)
with the same transformation properties valid for the covector components. With a slight abuse of notation, we use $t$ to denote the time component and $a$ for the spatial components of the spacetime index. Accordingly, the electromagnetic field strength tensor transforms as
$\mathds{P}F^{ta}=$ $=\mathds{P}\left(\nabla^{t}A^{a}-\nabla^{a}A^{t}\right)=-F^{ta}\,,$ (139)
$\mathds{P}F^{ab}=$ $=\mathds{P}\left(\nabla^{a}A^{b}-\nabla^{b}A^{a}\right)=F^{ab}\,.$ (140)
Finally, the conjugate momentum $\pi^{a}$ associated with the spatial components of the electromagnetic 4-vector potential transforms as
$\mathds{P}\pi^{a}=-\pi^{a}\,.$ (141)
Having established how the various variables transform under parity, we are able to study the behavior of the symplectic term and the constraints of the system. The symplectic term can be shown to transforms as
$\begin{split}\mathds{P}\left(\frac{1}{\kappa\beta}\int d^{4}xE_{i}^{a}\mathcal{L}_{t}A_{a}^{i}\right)=&\frac{1}{\kappa\beta}\int d^{4}x(-E_{i}^{a})\mathcal{L}_{t}\left(\tilde{\Gamma}_{a}^{i}-\beta\left(\tilde{K}_{a}^{i}-\frac{\kappa}{4}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\right)\\ =&\frac{1}{\kappa\beta}\int d^{4}xE_{i}^{a}\mathcal{L}_{t}\left(\beta\left(\tilde{K}_{a}^{i}-\frac{\kappa}{4}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\right)\\ =&\frac{1}{\kappa\beta}\int d^{4}xE_{i}^{a}\mathcal{L}_{t}A_{a}^{i}\,.\end{split}$ (142)
In deriving this result, we used the fact that $\tilde{\Gamma}_{a}^{i}$ is a function of the triad and, consequently, has a vanishing Poisson bracket with the densitized triad, i.e., $\{E_{i}^{a},\tilde{\Gamma}_{b}^{j}\}=0$. This result implies that the symplectic term is invariant under parity transformation.
For the Gauss constraint, we do not need to analyse the case in which the Lagrange multiplier has internal indices, since, as mentioned before, we are performing our analysis on the hypersurface in which $\mathcal{G}_{i}^{\text{G+F}}=0$. For the terms in the Gauss constraint that depend only on the spacetime indices, we obtain
$\mathcal{G}^{\text{QED}}=\nabla_{a}\pi^{a}-\sqrt{\text{det}\,h}\,q\mathcal{J}^{0}\,.$ (143)
Using the transformation rules for spacetime indices, we can conclude that this part of the constraint is also invariant under parity.
Applying the parity operator to the diffeomorphism constraint and considering again the hypersurface $\mathcal{G}_{i}^{\text{G+F}}=0$, one obtains
$\mathds{P}\left(\mathcal{H}_{a}^{\text{\text{Full}}}\right)=$ $\mathcal{H}_{a}^{\text{\text{Full}}}-\frac{2}{\kappa\beta}\Bigl{[}E_{i}^{b}\partial_{a}\tilde{\Gamma}_{b}^{i}+\tilde{\Gamma}_{a}^{i}\partial_{b}E_{i}^{b}+\epsilon^{i}{}_{jk}\tilde{\Gamma}_{a}^{j}\tilde{\Gamma}_{b}^{k}E_{i}^{b}$
$+\beta^{2}\epsilon^{i}{}_{jk}\left(\tilde{K}_{a}^{j}-\frac{\kappa}{4}\epsilon^{j}{}_{mn}e_{a}^{m}J^{n}\right)\left(\tilde{K}_{b}^{k}-\frac{\kappa}{4}\epsilon^{k}{}_{pq}e_{b}^{p}J^{q}\right)\sqrt{\text{det}\,h}e_{i}^{b}\Bigr{]}-\sqrt{\text{det}\,h}\beta\tilde{K}_{a}^{i}J_{i}$
$=$ $\mathcal{H}_{a}^{\text{\text{Full}}}-\frac{1}{\kappa\beta}\Bigl{[}2E_{i}^{b}\partial_{a}\tilde{\Gamma}_{b}^{i}+\beta^{2}\epsilon^{i}{}_{jk}\left(\tilde{K}_{a}^{j}-\frac{\kappa}{4}\epsilon^{j}{}_{mn}e_{a}^{m}J^{n}\right)(-\kappa J_{j}\sqrt{\text{det}\,h})\Bigr{]}$
$-\sqrt{\text{det}\,h}\beta\tilde{K}_{a}^{i}J_{i}$
$=$ $\mathcal{H}_{a}^{\text{\text{Full}}}-\frac{2}{\kappa\beta}E_{i}^{b}\partial_{a}\tilde{\Gamma}_{b}^{i}$
$=$ $\mathcal{H}_{a}^{\text{\text{Full}}}-\frac{2}{\kappa\beta}\epsilon_{ij}{}^{k}\tilde{\Gamma}_{b}^{i}\tilde{\Gamma}_{a}^{j}E_{k}^{b}$
$=$ $\mathcal{H}_{a}^{\text{\text{Full}}}\,.$ (144)
In the above derivation, we have used the fact that $\epsilon_{ij}{}^{k}E_{k}^{a}\tilde{\Gamma}_{a}^{i}=0$ with $\tilde{\Gamma}_{a}^{i}$ written in terms of the triad.
Finally, we examine the behavior of the Hamiltonian constraint under a parity transformation. All previously established comments and transformation rules remain valid. Therefore, after a series of straightforward yet lengthy calculations, we conclude that the Hamiltonian constraint is also invariant under parity transformation, i.e.,
$\mathds{P}\mathcal{H}^{\text{\text{Full}}}=$ $\mathcal{H}^{\text{\text{Full}}}+\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(-4\mathcal{D}_{a}^{(\tilde{\Gamma})}\tilde{K}_{b}^{k}+\beta\kappa\mathcal{D}_{a}^{(\tilde{\Gamma})}(\epsilon^{k}{}_{pq}e_{b}^{p}J^{q})\right)+\beta E_{i}^{a}\mathcal{D}_{a}^{(\tilde{\Gamma})}J^{i}$
$=$ $\mathcal{H}^{\text{\text{Full}}}-\frac{2}{\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\mathcal{D}_{a}^{(\tilde{\Gamma})}\tilde{K}_{b}^{k}$
$=$ $\mathcal{H}^{\text{\text{Full}}}$ (145)
As a consequence of the analysis in this subsection, we conclude that the theory demonstrates full invariance under parity transformations, which serves as another consistency check of our formulation.
4.3 Minimal vs. non-minimal coupling
So far, our analysis has been based on the assumption of minimal coupling between gravity and fermionic fields. However, some authors have pointed out that this procedure may not be adequate in curved spacetime Bojowald:2010qpa ; PhysRevD.73.084016 ; Alexandrov:2008iy . In particular, they found that employing minimal coupling can result in a system that lacks parity invariance. As noted in Das:2008r , this issue may stem from sign inconsistencies or the choice of representation for the gamma matrices. Although, in our work we did not encounter this problem, nevertheless, for the sake of completeness, we now examine how our earlier results are affected when non-minimal coupling is assumed, focusing specifically on the case of Weyl fermions. Generalization to Dirac spinors is straightforward.
To motivate the need for this modification, we begin by highlighting a potentially problematic term in the minimally coupled gravity-fermion action:
$S_{\text{G+F}}=\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}F_{ab}^{KL}(\omega)+\frac{i}{2}\int d^{4}x|e|\Bigl{\{}e_{I}^{a}\overline{\Psi}\gamma^{I}\mathfrak{D}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}\gamma^{I}\Psi\Bigr{\}}\,.$ (146)
As pointed out before, the presence of fermions modifies the compatibility condition. The difference between the full covariant derivative and the torsion-free one can be expressed as
$(\mathcal{D}_{a}-\tilde{\mathcal{D}}_{a})X_{I}=C_{aI}\,^{J}X_{J}$ (147)
where $C_{a}^{IJ}$ is the contorsion tensor. The corresponding curvature then becomes
$F_{ab}^{IJ}=\tilde{F}_{ab}^{IJ}+2\partial_{[a}C_{b]}^{IJ}+[C_{a},C_{b}]^{IJ}$ (148)
Substituting this expression into the action $S_{\text{G+F}}$, we obtain
$\begin{split}S_{\text{G+F}}=&\tilde{S}_{\text{G}}+\tilde{S}_{\text{F}}+S_{\text{int}}\\ =&\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}\tilde{F}_{ab}^{KL}(\tilde{\omega})+\frac{i}{2}\int d^{4}x|e|\Bigl{\{}e_{I}^{a}\overline{\Psi}\gamma^{I}\tilde{\mathfrak{D}}_{a}\Psi-e_{I}^{a}\overline{\tilde{\mathfrak{D}}_{a}\Psi}\gamma^{I}\Psi\Bigr{\}}\\ &+\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}[C_{a},C_{b}]^{KL}-\frac{1}{4}\int d^{4}x|\text{det}\,e|\epsilon^{I}\,_{JKL}e^{a}_{I}C_{a}^{JK}J^{L}\,.\end{split}$ (149)
Here, $\tilde{S}_{\text{G}}$ and $\tilde{S}_{\text{F}}$ are respectively the gravitational and fermionic actions constructed using the torsion-free spin connection 1-form $\tilde{\omega}_{a}^{IJ}$ and $S_{\text{int}}$ is the minimal coupling interaction term between gravity and fermions. After performing the contractions in the interaction terms, the full action becomes:
$S_{\text{G+F}}=\tilde{S}_{\text{G}}+\tilde{S}_{\text{F}}-\frac{3\kappa}{16}\frac{\beta^{2}}{1+\beta^{2}}\int d^{4}x|\text{det}\,e|J_{I}J^{I}.$ (150)
This coupling term explicitly depends on the Barbero-Immirzi parameter $\beta$, which some authors believe would be problematic PhysRevD.73.084016 . It is thus suggested that non-minimal couplings could provide a more consistent framework since setting the coupling constant $\alpha=-\frac{1}{\beta}$ yields a $\beta$-independent interaction term. Hence, we present the idea behind this suggestion and the corresponding modifications to the constraints and compatibility conditions when a non-minimal interaction is introduced.
The action incorporating a non-minimal coupling of gravity to Weyl fermions can be written as
$\begin{split}S_{\text{G+F}}=&\frac{1}{2\kappa}\int d^{4}x|\text{det}\,e|e_{I}^{a}e_{J}^{b}P_{KL}^{IJ}F_{ab}^{KL}(\omega)\\ &+\frac{i}{2}\int d^{4}x|e|\Bigl{\{}e_{I}^{a}\overline{\Psi}\gamma^{I}(1-i\alpha\gamma_{5})\mathfrak{D}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}(1-i\bar{\alpha}\gamma_{5})\gamma^{I}\Psi\Bigr{\}}\,,\end{split}$ (151)
where $\alpha\in\mathbb{C}$ parametrizes the non-minimal coupling. Following the same procedure as discussed in detail in previous sections, the constraints become
$\mathcal{G}_{i}^{\text{G+F}}=$ $\frac{1}{\kappa\beta}\mathcal{D}_{a}^{(A)}E_{i}^{a}+\frac{1}{2}\sqrt{\text{det}\,h}J_{i}+\frac{i}{2}(\alpha-\bar{\alpha})\sqrt{\text{det}\,h}\mathcal{J}_{i},$ (152)
$\mathcal{S}_{i}^{\text{G+F}}=$ $-\frac{1+\beta^{2}}{\kappa\beta}\epsilon_{ij}{}^{k}K_{a}^{j}E_{k}^{a}-\frac{1}{2}\beta\sqrt{\text{det}\,h}\,J_{i}-\frac{1}{4}(-\alpha+\bar{\alpha})\sqrt{\text{det}\,h}J_{i}$
$-\frac{i\beta}{2}(\alpha-\bar{\alpha})\sqrt{\text{det}\,h}\mathcal{J}_{i},$ (153)
$\mathcal{H}^{\text{G+F}}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\Bigl{\{}\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\Bigr{\}}$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}(1-i\alpha\gamma_{5})\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}(1-i\bar{\alpha}\gamma_{5})\gamma^{i}\Psi\right),$ (154)
$\mathcal{H}_{a}^{\text{G+F}}=$ $\frac{1}{\kappa\beta}\left(E_{j}^{b}\mathcal{F}_{ab}^{j}(A)-(1+\beta^{2})\epsilon^{j}{}_{kl}E_{j}^{b}K_{a}^{k}K_{b}^{l}\right)$
$-\frac{i}{2}\sqrt{\text{det}\,h}\left(\overline{\Psi}\gamma^{0}(1-i\alpha\gamma_{5})\mathfrak{D}_{a}\Psi+\overline{\mathfrak{D}_{a}\Psi}(1-i\bar{\alpha}\gamma_{5})\gamma^{0}\Psi\right).$ (155)
In particular case of a real coupling $\alpha=\bar{\alpha}$, the constraints simplify and the new compatibility condition becomes
$\begin{split}\mathcal{D}_{b}(|e|e_{K}^{[a}e_{L}^{b]})=&\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}|e|\Bigl{[}\left(1-\frac{\alpha}{\beta}\right)\epsilon_{KL}{}^{MN}e_{M}^{a}J_{N}-\left(\frac{1}{\beta}+\alpha\right)\left(e_{K}^{a}J_{L}-e_{L}^{a}J_{K}\right)\Bigr{]}.\end{split}$ (156)
Consequently, the spin connection is modified as follow
$\Gamma_{a}^{i}=\tilde{\Gamma}_{a}^{i}+\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\Bigl{[}\left(1-\frac{\alpha}{\beta}\right)e_{a}^{i}J^{0}-\left(\frac{1}{\beta}+\alpha\right)\epsilon^{i}\,_{jk}e_{a}^{j}J^{k}\Bigr{]}.$ (157)
Finally, imposing the new compatibility condition on the Gauss constraint leads to
$\mathcal{G}_{i}^{\text{G+F}}=\frac{1}{\kappa}\epsilon_{ij}\,^{k}K_{a}^{j}E_{k}^{a}+\frac{1}{2}\frac{\beta^{2}}{1+\beta^{2}}\left(1-\frac{\alpha}{\beta}\right)\sqrt{\text{det}\,h}J_{i}$ (158)
5 Discussion and conclusions
In this work we have provided the full analysis of the canonical formulation of the system of gravity, fermions, and photons, and their interaction in a background-independent way. In particular, we have considered the interaction term of the QED, which to our knowledge, has not been considered in previous studies.
We have analyzed the constraint structure of the system in the Ashtekar-Barbero variables, with either Weyl or Dirac fermions, and have shown how the connection variables are modified due to the presence of fermions. Furthermore, the expression for the Hamiltonian, spatial diffeomorphism and Guass constraints are given in a way that the contribution of each of the sectors, i.e., gravity, fermions, and photons, are clear. We have also derived the EOM of the matter sector, which reduce to the standard EOM of QFT on fixed curved or flat spacetime in the appropriate limits. The behavior of the system under parity transformation is studied in detail. Finally, we presented the modifications to the constraints arising from the introduction of non-minimal coupling between gravity and fermions.
This analysis and the results pave the way for further canonical analysis of realistic matter in the universe involved in electromagnetic interactions in strong and dynamical (non-fixed) gravity regimes. Most importantly, this work opens the door for further detailed studies of an interactive system (QED in a dynamical background) using loop quantum gravity, polymer quantization, generalized uncertainty principle, or other theory that uses canonical quantization techniques. A rather immediate application of this analysis is in symmetry-reduced systems such as cosmological models involving electromagnetic matter, or collapse scenarios in spherical symmetric or rotating black holes, which up to now, have been done almost exclusively only with scalar fields.
Acknowledgements.
The authors acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). We also would like to thank Mehdi Assanioussi and Klaus Liegener for insightful discussions.
Appendix A Complementary information on fermions
Appendix A serves to clarify the notation adopted for the fermionic field, the gamma matrices, and their associated identities. Given the ambiguity that exists in the literature, this appendix establishes the conventions used throughout the paper to ensure consistency.
A.1 Notes on helicity and chirality
In this subsection, we introduce some key concepts related to helicity and chirality Pal:2011pb ; Schwartz:2014sze , which will help clarify the discussions that follow.
A general definition of “handedness” applicable to any particle involves the relative orientation between its momentum and angular momentum. For a particle with three-momentum $\mathbf{p}$, helicity is defined as
$h_{\mathbf{p}}\coloneqq\frac{2\mathbf{J}\cdot\mathbf{p}}{|\mathbf{p}|}\,,$ (159)
where $\mathbf{J}$ is the total angular momentum of the particle. Since the orbital angular momentum is always perpendicular to the direction of motion, it does not contribute to helicity. Therefore, helicity effectively measures twice the projection of the spin along the direction of momentum. Defining the spin matrices as $\frac{1}{2}\Sigma$, the helicity operator can be rewritten as
$h_{\mathbf{p}}\coloneqq\frac{\Sigma\cdot\mathbf{p}}{|\mathbf{p}|}\,.$ (160)
It follows that the eigenvalues of $h_{\mathbf{p}}$ are ±$\pm 1$. An eigenstate with eigenvalue $-1$ is referred to as “left-handed”, while one with eigenvalue $+1$ is called “right-handed”.
A key property of helicity is that it commutes with the Dirac Hamiltonian, which means it is conserved in coordinate time. Moreover, helicity is invariant under rotations, as evident from the rotational invariance of the dot product in its definition. In other words, different observers related by a spatial rotation will agree on the helicity of a given particle. However, helicity is not invariant under Lorentz boosts. As a result, a massive fermion [^3] The change in helicity can only occur in frames moving faster than the particle. Since such a frame cannot exist for a massless particle, helicity remains invariant in that case. cannot be strictly left- or right-helical in an absolute sense; the helicity depends on the observer’s frame of reference.
With regard to the definition of chirality, it is associated with the matrix $\gamma_{5}$ (see Appendix A.2 for its properties). This matrix allows us to construct projection operators that act on fermion fields as follows
$\mathds{L}=$ $\frac{1}{2}\left(1-\gamma_{5}\right),$ $\mathds{R}=$ $\frac{1}{2}\left(1+\gamma_{5}\right)\,.$ (161)
These projection operators are commonly referred to as “left-chiral” and “right-chiral”, respectively. Given any Dirac spinor $\Psi(x)$ that satisfies the Dirac equation, we can decompose it into its left- and right-chiral components as
$\Psi=\Psi_{L}+\Psi_{R},$ (162)
where
$\Psi_{L}=$ $\mathds{L}\Psi,$ $\Psi_{R}=$ $\mathds{R}\Psi\,.$ (163)
An important feature of chirality is that a left-chiral solution to the Dirac equation remains left-chiral under Lorentz transformations. This is ensured by the fact that $\gamma_{5}$ anticommutes with all gamma matrices, allowing the chiral projections to be defined in a Lorentz covariant manner. However, chirality is not conserved for a massive free particle, as $\gamma_{5}$ does not commute with the mass term in the Dirac Hamiltonian. Since the mass term involves only one gamma matrix, it anticommutes with $\gamma_{5}$, leading to a mixing of chiral components. In contrast, there is no such issue with the kinetic (derivative) term, which preserves chirality.
In this sense, chirality and helicity exhibit complementary behaviors: helicity is conserved for a free particle but is not Lorentz invariant, whereas chirality is Lorentz invariant but not conserved when the particle has mass.
In the massless case, applying the Dirac equation to a spinor eigenstate of 3-momentum $\mathbf{p}$, we find
$\gamma_{5}w_{\mathbf{p}}=\frac{\Sigma\cdot\mathbf{p}}{|\mathbf{p}|}w_{\mathbf{p}}\,,$ (164)
where $w_{\mathbf{p}}$ denotes a general spinor. This relation shows that chirality and helicity coincide for massless spinors: the action of $\gamma_{5}$ on the spinor yields its helicity. In this limit, chirality becomes a physical observable and Lorentz-invariant quantity.
When mass is introduced, the equations of motion cause mixing between left- and right-chiral components. However, since momentum and spin remain good quantum numbers in the free theory —even for massive particles— helicity is still conserved. As a result, helicity remains a useful and meaningful concept in the massive case. The crucial distinction is that, in the presence of mass, helicity eigenstates no longer coincide with chirality eigenstates.
A.2 Weyl fermions in flat spacetime
In order to maintain consistency with the majority of the literature regarding fermionic matter Schwartz:2014sze ; Weinberg:1995s ; Peskin:1995ev , we will switch to the standard QFT signature convention, (+)(+\,-\,-\,-)( + - - - ), for Section A.2 only.
In general, elements of the vector space on which the spin-$\frac{1}{2}$ representations act are known as spinors. When we consider massless fermions, we can talk of positive or negative helicity fermion or of left or right chiral fermion  [^4] In the massless case, helicity are chiraly are related to each other (see Appendix A.1).. It is well known that the proper Lorentz algebra $\mathfrak{so}(1,3)$ is isomorphic to the Lie algebra of $SU(2)\times SU(2)$, meaning that any representation of the Lorentz algebra can be characterized by how it transforms under each of the two $\mathfrak{su}(2)$ factors. In this framework, a left-chiral fermion transforms as a doublet under one $\mathfrak{su}(2)$ and as a singlet under the other, denoted by the representation $(\frac{1}{2},0)$. Conversely, a right-chiral fermion transforms as $(0,\frac{1}{2})$. These are known as Weyl fermions Pal:2011pb ; Schwartz:2014sze
$\psi_{R}=$ $\begin{pmatrix}\psi_{1}(x)\\ \psi_{2}(x)\end{pmatrix},$ for the (0,12) $\text{for the $\left(0,\frac{1}{2}\right)$ representation}\,,$ (165)
$\psi_{L}=$ $\begin{pmatrix}\psi_{1}(x)\\ \psi_{2}(x)\end{pmatrix},$ for the (12,0) $\text{for the $\left(\frac{1}{2},0\right)$ representation}\,.$ (166)
Under rotational angles $\theta_{i}$ and boost angles $\omega_{i}$, they transform as
$\psi_{L}\rightarrow\Lambda_{L}\psi_{L}=$ $e^{i\frac{\bar{\sigma}}{2}\left(\bar{\theta}+i\bar{\omega}\right)}\psi_{L},$ $\psi_{R}\rightarrow\Lambda_{R}\psi_{R}=$ $e^{i\frac{\bar{\sigma}}{2}\left(\bar{\theta}-i\bar{\omega}\right)}\psi_{R}\,.$ (167)
Since $\Lambda_{L}$ and $\Lambda_{R}$ are complex, so are $\psi_{L}$ and $\psi_{R}$ Similarly,
$\psi_{L}^{\dagger}\rightarrow$ $\psi_{L}^{\dagger}e^{-i\frac{\bar{\sigma}}{2}\left(\bar{\theta}-i\bar{\omega}\right)},$ $\psi_{R}^{\dagger}\rightarrow$ $\psi_{R}^{\dagger}e^{-i\frac{\bar{\sigma}}{2}\left(\bar{\theta}+i\bar{\omega}\right)}\,.$ (168)
By definition, Weyl fermions are massless, so we are only concerned with the kinetic term of their Lagrangian. To derive this term, consider the vector quantities $v^{i}=\psi_{R}^{\dagger}\sigma^{i}\psi_{R}$ and $\tilde{v}^{i}=\psi_{L}^{\dagger}\sigma^{i}\psi_{L}$, where $\sigma^{i}$ denotes the $i$th Pauli matrix. Using the well-known identities of the Pauli matrices,
$\left[\sigma^{i},\sigma^{j}\right]=$ $2i\epsilon^{ij}\,_{k}\sigma^{k}\,,$$\left\{\sigma^{i},\sigma^{j}\right\}=$ $2\delta^{ij}\mathbb{1}\,,$ (169)
we can compute the infinitesimal transformations of $v^{i}$ and $\tilde{v}^{i}$ under Lorentz transformations. This yields
$\delta v^{i}=$ $-\epsilon^{ij}\,_{k}\theta_{j}\psi_{R}^{\dagger}\sigma^{k}\psi_{R}+\omega^{i}\psi_{R}^{\dagger}\psi_{R}\,,$ $\delta\tilde{v}^{i}=$ $-\epsilon^{ij}\,_{k}\theta_{j}\psi_{L}^{\dagger}\sigma^{k}\psi_{L}-\omega^{i}\psi_{L}^{\dagger}\psi_{L}\,.$ (170)
Thus, both $v^{i}$ and $\tilde{v}^{i}$ transform as the spatial parts of a Lorentz 4-vector $\delta V^{\mu}=\Lambda_{\nu}^{\mu}V^{\nu}$ with $\Lambda_{i}^{0}=\omega^{i}$ and $\Lambda_{j}^{i}=\epsilon^{i}\,_{jk}\theta^{k}$. We can define
$v^{\mu}=$ $(\psi_{R}^{\dagger}\psi_{R},\psi_{R}^{\dagger}\sigma^{i}\psi_{R})\,,$ $\tilde{v}^{\mu}=$ $(\psi_{L}^{\dagger}\psi_{L},-\psi_{L}^{\dagger}\sigma^{i}\psi_{L})\,,$ (171)
so that the combinations $\psi_{R,L}^{\dagger}\partial_{t}\psi_{R,L}$ transform as Lorentz invariants.
Defining the matrices
$\sigma^{\mu}=$ $(\mathbb{1},\sigma^{i})\,,$ $\tilde{\sigma}^{\mu}=$ $(\mathbb{1},-\sigma^{i})\,,$ (172)
we can write the previous vectors in the following form
$v^{\mu}=$ $\psi_{R}^{\dagger}\sigma^{\mu}\psi_{R}\,,$ $\tilde{v}^{\mu}=$ $\psi_{L}^{\dagger}\tilde{\sigma}^{\mu}\psi_{L}\,,$ (173)
and the kinetic terms of the Lagrangian become
$\mathcal{L}_{R}=$ $i\psi_{R}^{\dagger}\sigma^{\mu}\partial_{\mu}\psi_{R}\,,$ $\mathcal{L}_{L}=$ $i\psi_{L}^{\dagger}\tilde{\sigma}^{\mu}\partial_{\mu}\psi_{L}\,.$ (174)
Hence, the Dirac Lagrangian for Weyl fermions can be expressed as
$\mathcal{L}_{\text{Dirac}}=i\psi_{R}^{\dagger}\sigma^{\mu}\partial_{\mu}\psi_{R}+i\psi_{L}^{\dagger}\tilde{\sigma}^{\mu}\partial_{\mu}\psi_{L}\,.$ (175)
It is now possible to introduce a notation that can be used for arbitrary fermionic field combining the two spinors into a four-component object
$\Psi=$ $\begin{pmatrix}\psi_{L}\\ \psi_{R}\end{pmatrix},$ $\Psi^{\dagger}=$ $\begin{pmatrix}\psi_{L}^{\dagger}&\psi_{R}^{\dagger}\end{pmatrix}.$ (176)
In this basis, known as the Weyl (or chiral) representation, the gamma matrices take the form
$\gamma^{\mu}=\begin{pmatrix}0&\sigma^{\mu}\\ \tilde{\sigma}^{\mu}&0\end{pmatrix}\,,$ (177)
with components
$\gamma^{0}=$ $\begin{pmatrix}0&\mathbb{1}\\ \mathbb{1}&0\end{pmatrix}\,,$ $\gamma^{i}=$ $\begin{pmatrix}0&\sigma^{i}\\ -\sigma^{i}&0\end{pmatrix}\,.$ (178)
These matrices satisfy the Clifford algebra $\{\gamma^{\mu},\gamma^{\nu}\}=2\eta^{\mu\nu}$.
Moreover, we need to identify a suitable way to combine fermionic fields in order to construct a Lorentz scalar. This is achieved by introducing the Dirac adjoint (or conjugate) of the fermionic field, defined using the zeroth gamma matrix
$\overline{\Psi}=\Psi^{\dagger}\gamma^{0}=\begin{pmatrix}\psi_{L}^{\dagger}&\psi_{R}^{\dagger}\end{pmatrix}\begin{pmatrix}0&\mathbb{1}\\ \mathbb{1}&0\end{pmatrix}=\begin{pmatrix}\psi_{R}^{\dagger}&\psi_{L}^{\dagger}\end{pmatrix}\,.$ (179)
This definition ensures that the bilinear form $\overline{\Psi}\Psi$ transforms as a Lorentz scalar, and that the kinetic term $\overline{\Psi}\gamma^{\mu}\partial_{\mu}\Psi$ transforms as a Lorentz scalar density, preserving the overall covariance of the Lagrangian. Thus, we arrive at the conventional form of the Dirac Lagrangian for a Weyl fermionic field
$\mathcal{L}_{\text{Dirac}}=i\overline{\Psi}\gamma^{\mu}\partial_{\mu}\Psi\,.$ (180)
A.3 Weyl fermions in curved spacetime
Consider the fermionic field in a curved spacetime. In this context, it is necessary to introduce a suitable covariant derivative that accounts for the coupling between fermions and gravity. This is achieved by incorporating the generalized connection 1-form $\omega_{a}^{IJ}$. Hence, The covariant derivatives acting on the spinor field $\Psi$ and its conjugate field $\overline{\Psi}$ are defined respectively as
$\mathfrak{D}_{\mu}\Psi=$ $\partial_{\mu}\Psi+\frac{1}{2}\omega_{\mu}^{IJ}\sigma_{IJ}\Psi\,,$ (181)
$\overline{\mathfrak{D}_{\mu}\Psi}$ $=\partial_{\mu}\overline{\Psi}+\frac{1}{2}\omega_{\mu}^{IJ}\Psi^{\dagger}\sigma_{IJ}^{\dagger}\gamma^{0}\,,$ (182)
where $\sigma^{\mu\nu}\coloneqq\frac{1}{4}[\gamma^{\mu},\gamma^{\nu}]$ are the generators of the Lorentz group in the spinor representation. Using this formalism, the action for a Weyl fermion in curved spacetime is given by
$S_{F}=\frac{i}{2}\int d^{4}x|e|\Bigl{\{}e_{I}^{\mu}\overline{\Psi}\gamma^{I}\mathfrak{D}_{\mu}\Psi-e_{I}^{\mu}\overline{\mathfrak{D}_{\mu}\Psi}\gamma^{I}\Psi\Bigr{\}}\,.$ (183)
However, in General Relativity we adopt the mostly-plus metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ), which requires a modification of some algebraic relations involving the gamma matrices. While the Pauli matrices and their commutation and anticommutation relations remain unchanged, the Clifford algebra for the gamma matrices must be adapted accordingly. We choose to retain the explicit form of the gamma matrices as given in Eq.(178). However, due to the change in metric signature, the anticommutation relation now reads:
$\left\{\gamma^{I},\gamma^{J}\right\}=-2\eta^{IJ}\,,$ (184)
where $\eta^{IJ}$ is the Minkoswki metric. This relation leads to the following properties for the gamma matrices:
$\left(\gamma^{0}\right)^{\dagger}=\gamma^{0}$ $\implies\left(\gamma^{0}\right)^{2}=\mathbb{1}\,,$ $\left(\gamma^{i}\right)^{\dagger}=-\gamma^{i}$ $\implies\left(\gamma^{i}\right)^{2}=-\mathbb{1}\,,$ (185)
which imply that $\gamma^{0}$ is Hermitian and unitary, while $\gamma^{i}$ are anti-Hermitian and unitary.
Moreover, adopting the GR signature, the definition of the $\gamma^{5}$ matrix becomes
$\gamma^{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}=\begin{pmatrix}-\mathbb{1}&0\\ 0&\mathbb{1}\end{pmatrix}\,.$ (186)
Unlike other approaches found in the literature Bojowald:2010qpa ; Das:2008r , we do not modify the gamma matrices or the fermionic fields when transitioning to the curved spacetime formulation of the Dirac action. The key differences between the flat and curved spacetime cases arise from the introduction of the tetrad fields and a sign change in the anticommutation relations of the gamma matrices, as shown in Eq. (184). As a consequence, several familiar identities involving gamma matrices are altered. Below, we present a list of these modified identities, which can be verified straightforwardly
$\gamma^{I}\gamma_{I}=$ $-4\mathbb{1}\,,$ (187)
$\gamma^{I}\gamma^{J}\gamma_{I}=$ $2\gamma^{I}\,,$ (188)
$\gamma^{I}\gamma^{J}\gamma^{K}\gamma_{I}=$ $4\eta^{JK}\mathbb{1}\,,$ (189)
$\gamma^{I}\gamma^{J}\gamma^{K}\gamma^{L}\gamma_{I}=$ $2\gamma^{L}\gamma^{K}\gamma^{J}\,,$ (190)
$\gamma^{I}\gamma^{J}\gamma^{K}=$ $-\eta^{IJ}\gamma^{K}-\eta^{JK}\gamma^{I}+\eta^{IK}\gamma^{J}+i\epsilon^{LIJK}\gamma_{L}\gamma^{5}\,,$ (191)
$\epsilon^{IJKL}\sigma_{IJ}=$ $-2i\gamma^{5}\sigma^{KL}\,,$ (192)
$\gamma^{I}\sigma_{JK}+\sigma_{JK}\gamma^{I}=$ $\,i\epsilon^{I}\,_{JKL}\gamma_{5}\gamma^{L}\,,$ (193)
$\sigma_{JK}\gamma^{I}-\gamma^{I}\sigma_{JK}$ $=\eta_{J}^{I}\gamma_{K}-\eta_{K}^{I}\gamma_{J}\,.$ (194)
A.4 Dirac fermions in flat and curved spacetime
A general fermion field transforms according to a reducible representation [^5] A representation is called reducible if it can be decomposed into a direct sum of irreducible representations, each of which transforms independently under the group. $(\frac{1}{2},0)\oplus(0,\frac{1}{2})$. This means that it can be described in terms of two irreducible Weyl components: one left-chiral and one right-chiral. This decomposition underscores the fundamental role of Weyl fields, which serve as the basic building blocks from which any fermionic field can be constructed.
We begin by considering the case of flat spacetime. To construct a massive fermionic field, we consider the Lorentz-invariant combinations $\psi_{L}^{\dagger}\psi_{R}$ and $\psi_{R}^{\dagger}\psi_{L}$, which can be used to introduce a mass term in the Dirac Lagrangian:
$\mathcal{L}_{\text{mass}}=-m(\psi_{L}^{\dagger}\psi_{R}+\psi_{R}^{\dagger}\psi_{L})\,.$ (195)
When combined with the kinetic term, the resulting theory can describe Majorana fermions, provided the fields are real solutions of the Dirac equation. This would require imaginary gamma matrices. However, this scenario will not be explored further in this work. Instead, we will focus on generic complex solutions of the Dirac equation, without imposing any reality conditions. This leads us to the Dirac field, whose Lagrangian in terms of the Weyl components is given by
$\mathcal{L}_{\text{Dirac}}=i\psi_{R}^{\dagger}\sigma^{\mu}\partial_{\mu}\psi_{R}+i\psi_{L}^{\dagger}\tilde{\sigma}^{\mu}\partial_{\mu}\psi_{L}-m(\psi_{L}^{\dagger}\psi_{R}+\psi_{R}^{\dagger}\psi_{L})\,.$ (196)
The Dirac field is particularly useful for studying QED, which does not distinguish between left- and right-handed components, i.e., it is not a chiral theory. Since we are not concerned with analyzing chirality in our discussion, working in the standard (Dirac) representation is more convenient.
One of the future goals of this study will be to examine propagators in a modified version of QED. Importantly, the choice of representation for the gamma matrices (whether Dirac, Weyl, or Majorana) does not affect the physical structure or form of the propagators. Therefore, when referring to the Dirac field, we will adopt the four-component spinor formalism as defined in (176), and in this case, the Dirac Lagrangian takes the form
$\mathcal{L}_{\text{Dirac}}=i\overline{\Psi}\gamma^{\mu}\partial_{\mu}\Psi-m\overline{\Psi}\Psi\,,$ (197)
where the gamma matrices are given in Eq.(178).
To extend the previously discussed Lagrangian to curved spacetime, we follow the procedure outlined in Appendix A.3. The fermionic covariant derivative is introduced as defined in Eqs. (181)-(182). It is important to note that the mass term remains unchanged in curved spacetime, as it does not involve any explicit gamma matrix or derivative operators. As a result, the action for Dirac fields in a curved background takes the form
$S_{F}=\int d^{4}x|e|\left\{\frac{i}{2}\left(e_{I}^{a}\overline{\Psi}\gamma^{I}\mathfrak{D}_{a}\Psi-e_{I}^{a}\overline{\mathfrak{D}_{a}\Psi}\gamma^{I}\Psi\right)-m\overline{\Psi}\Psi\right\}\,.$ (198)
where all symbols carry their usual meaning.